### v2

# Preliminaries on Pre-Hilbert Structures on Polynomial Spaces and Associated Laplacians

Abstract.
We study orthogonal polynomial systems arising from general pre-Hilbert inner products on polynomial spaces, beyond the classical framework of measures. To each such inner product we associate a canonical Laplacian defined from an abstract derivation, and we investigate the operator-theoretic structures induced by this construction.
Our main contribution is the introduction of a resolvent-based distance between polynomial Hilbert geometries, and the proof of quantitative stability results for finite-degree orthogonalization procedures. In particular, we show that norm-resolvent closeness of the associated Laplacians implies stability of Gram–Schmidt orthogonal bases, orthogonal projectors and reproducing kernels on all finite-dimensional polynomial subspaces.
The general theory is illustrated by several explicit examples. We analyze in detail the case of orthogonal polynomials on the unit circle, comparing classical $L^{2}$ geometries associated with finite Radon measures and Sobolev-type regularizations via Fourier methods. We also revisit the thin annulus problem, showing that its asymptotic regime admits a natural interpretation as a resolvent limit of polynomial geometries.
These results provide a unified operator-theoretic framework for the study of stability, degenerations and geometric limits of orthogonal polynomial systems.
Key words and phrases:
Orthogonal polynomials, Sobolev inner products, Gram–Schmidt stability, Laplacian operators, resolvent convergence, polynomial Hilbert geometries
2020 Mathematics Subject Classification:
42C05, 47A55, 46E35, 47B36
1. Introduction
Orthogonal polynomials play a central role in analysis, approximation theory and spectral theory. Classically, they are defined through $L^{2}$ inner products associated with measures, leading to well-known families such as Jacobi, Laguerre, Hermite or orthogonal polynomials on the unit circle see, e.g., [23, 3, 7, 9, 21, 22]. In this setting, the algebraic, analytic and spectral properties of the polynomial systems are tightly linked to the underlying measure.
In many situations, however, orthogonal polynomials arise from inner products that are not purely measure-based. Sobolev and fractional Sobolev inner products, involving derivatives or nonlocal energies, provide natural examples motivated by approximation theory, partial differential equations and numerical analysis. More generally, one may consider arbitrary pre-Hilbert inner products on polynomial spaces, possibly defined by quadratic forms rather than measures. In such cases, classical tools based on moment problems or three-term recurrence relations are no longer sufficient to describe the underlying structure, as already observed for Sobolev and nonlocal inner products [15, 6]..
The purpose of this paper is to study orthogonal polynomial systems from an operator-theoretic and geometric perspective. Given a polynomial space equipped with a pre-Hilbert inner product and a derivation $D,$ we associate a canonical Laplacian operator defined as $\Delta=D^{*}D$. This operator directly depends on the interaction between the algebraic structure of polynomials and on the chosen inner product. Our approach emphasizes the role of the Laplacian as the fundamental object governing orthogonalization procedures. A central question addressed here is the stability of orthogonal polynomial systems under perturbations of the underlying inner product. Our key tool remains on the Laplacian $\Delta$. Indeed, rather than comparing measures or coefficients directly, we introduce a distance between polynomial Hilbert geometries based on the norm-resolvent difference of the associated Laplacians. Our distance is based on the norm of the resolvent difference $\|(1+\Delta_{1})^{-1}-(1+\Delta_{2})^{-1}\|$. Norm-resolvent convergence is classical in perturbation theory and spectral analysis, see, e.g., [8, 20, 4]. Related comparison topologies for (typically unbounded) operators include the gap and graph metrics [8, 12], as well as variational notions such as Mosco convergence for closed forms [17, 11]. What is specific to the present work is the use of a resolvent-based metric in the context of polynomial Hilbert geometries, in order to obtain quantitative stability of finite-degree Gram–Schmidt orthogonalization, projectors and reproducing kernels. Our main results show that resolvent closeness of Laplacians implies quantitative stability of Gram–Schmidt orthogonalization on all finite-dimensional polynomial subspaces. In particular, orthogonal projectors, orthonormal bases and reproducing kernels depend on the geometry in the resolvent topology. These results hold in a fully abstract framework and do not rely on the existence of an underlying measure. We use the norm of the resolvent difference as a metric on polynomial Hilbert geometries.
The abstract theory is complemented by explicit examples. We first analyze orthogonal polynomials on the unit circle, comparing classical $L^{2}$ geometries associated with finite Radon measures and Sobolev-type regularizations. Using Fourier methods, we derive detailed resolvent estimates and obtain quantitative stability results for orthogonal polynomials on the unit circle at fixed degree. We then revisit the thin annulus problem, showing that the collapse of a two-dimensional Sobolev geometry onto a one-dimensional limit admits a natural interpretation as a resolvent convergence of polynomial Laplacians.
Altogether, the results of this paper provide a unified framework for stability, degenerations and geometric limits of orthogonal polynomial systems, extending classical measure-based theory [23, 3, 7, 21, 22] to broader polynomial inner products such as Sobolev-type geometries [6, 15], in a spirit reminiscent of operator and form convergence in spectral theory [8, 17, 11].
2. Preliminaries
2.1. Polynomial spaces and pre-Hilbert structures
Let $\Omega\subset\mathbb{R}^{d}$ be a domain (i.e. an arbitrary subset, not necessarily open). We denote by
$\mathcal{P}(\Omega):=\{\,p|_{\Omega}:\ p\in\mathbb{C}[x_{1},\dots,x_{d}]\,\}$
the space of polynomial functions on $\Omega$. Let
$\langle\cdot,\cdot\rangle:\mathcal{P}(\Omega)\times\mathcal{P}(\Omega)\to\mathbb{C}$
be a positive semidefinite sesquilinear form (conjugate-linear in the first variable). Set
$N:=\{p\in\mathcal{P}(\Omega):\langle p,p\rangle=0\},\qquad\mathcal{H}_{0}:=\mathcal{P}(\Omega)/N.$
Then $\langle\cdot,\cdot\rangle$ descends to a genuine Hermitian inner product on $\mathcal{H}_{0}$, and we denote by $H$ the Hilbert completion of $(\mathcal{H}_{0},\langle\cdot,\cdot\rangle)$. We will systematically identify $\mathcal{H}_{0}$ with its canonical dense image in $H$.
Typical examples include:
  • $L^{2}$ inner products induced by finite Borel measures $\mu$ on $\Omega$,
    $\langle p,q\rangle_{L^{2}(\mu)}=\int_{\Omega}p(x)\,\overline{q(x)}\,d\mu(x),$
    as in the classical theory of orthogonal polynomials [23, 3, 21, 22];
  • Sobolev or fractional Sobolev inner products involving derivatives or nonlocal energies, see e.g. [14, 16];
  • discrete inner products arising from finite point sets and interpolation problems, [1, 24].
2.2. Gram–Schmidt orthogonalization and polynomial projectors
Fix an ordered family $(E_{n})_{n\geq 0}$ in $\mathcal{P}(\Omega)$ compatible with the degree filtration, for instance an enumeration of monomials with nondecreasing total degree. Assume that for each $N$ the family $(E_{0},\dots,E_{N})$ is linearly independent in $\mathcal{H}_{0}$. Define inductively the orthogonalized vectors $(\widetilde{p}_{n})_{n\geq 0}$ by
(2.1) $\widetilde{p}_{0}:=E_{0},\qquad\widetilde{p}_{n}:=E_{n}-\sum_{k=0}^{n-1}\frac{\langle E_{n},\widetilde{p}_{k}\rangle}{\langle\widetilde{p}_{k},\widetilde{p}_{k}\rangle}\,\widetilde{p}_{k},\quad n\geq 1,$
and the orthonormal polynomials $(p_{n})_{n\geq 0}$ by
(2.2) $p_{n}:=\frac{\widetilde{p}_{n}}{\|\widetilde{p}_{n}\|},\qquad\|\widetilde{p}_{n}\|^{2}=\langle\widetilde{p}_{n},\widetilde{p}_{n}\rangle.$
For each $N$, the subspace
$H_{\leq N}:=\mathrm{span}\{p_{0},\dots,p_{N}\}\subset H$
coincides with the image in $H$ of the polynomials of degree $\leq N$ (modulo $N$), and is finite-dimensional.
Gram matrices. Let $G_{N}\in\mathbb{C}^{(N+1)\times(N+1)}$ be the Gram matrix of $(E_{0},\dots,E_{N})$,
(2.3) $(G_{N})_{ij}:=\langle E_{j},E_{i}\rangle,\qquad 0\leq i,j\leq N.$
Then $G_{N}$ is Hermitian positive definite under the independence assumption above. Moreover, the coefficients in (2.1) can be written in matrix form: writing $E=(E_{0},\dots,E_{N})^{\top}$ and $\widetilde{p}=(\widetilde{p}_{0},\dots,\widetilde{p}_{N})^{\top}$, there exists a unique lower triangular matrix $L_{N}$ with positive diagonal such that
(2.4) $G_{N}=L_{N}L_{N}^{*},\qquad\widetilde{p}=L_{N}^{-1}E,$
and the normalized basis is $p=D_{N}^{-1}\widetilde{p}$ where $D_{N}=\mathrm{diag}(\|\widetilde{p}_{0}\|,\dots,\|\widetilde{p}_{N}\|)$.
Orthogonal projector. Let $P_{N}:H\to H_{\leq N}$ denote the orthogonal projector onto $H_{\leq N}$. In terms of the orthonormal basis $(p_{0},\dots,p_{N})$, one has for $f\in H$,
(2.5) $P_{N}f=\sum_{k=0}^{N}\langle f,p_{k}\rangle\,p_{k}.$
Equivalently, if we expand $P_{N}f=\sum_{j=0}^{N}c_{j}E_{j}$ in the (non-orthonormal) basis $(E_{0},\dots,E_{N})$, then the coefficient vector $c=(c_{0},\dots,c_{N})^{\top}$ is characterized by the normal equations
(2.6) $G_{N}\,c=b,\qquad b_{i}:=\langle f,E_{i}\rangle,\ 0\leq i\leq N,$
so $c=G_{N}^{-1}b$.
Finite-degree kernel (on compacts). For $x,y\in\Omega$, define the degree-$N$ kernel associated with $(p_{k})_{k=0}^{N}$ by
(2.7) $K_{N}(x,y):=\sum_{k=0}^{N}p_{k}(x)\,\overline{p_{k}(y)}.$
On a compact set $K\subset\Omega$, all polynomials are continuous, hence $K_{N}$ is continuous on $K\times K$. Moreover, $K_{N}$ is the matrix coefficient of $P_{N}$ evaluated on point-masses, in the sense made precise in Corollary 6.5 (kernel stability on compacts).
2.3. Derivations, closability, and associated Laplacians
Let $\widetilde{D}$ be a derivation on the polynomial algebra $\mathcal{P}(\Omega)$ (typically $\widetilde{D}$$\nabla$), taking values in $H^{m}$ (for some $m\geq 1$) and assume that
(2.8) $N\subset\ker\widetilde{D}$,
so that $\widetilde{D}$ induces a well-defined linear map $D_{0}:\mathcal{H}_{0}\to H^{m}$. Assume that $D_{0}$ is densely defined and closable. Denote by $D:=\overline{D_{0}}$ its closure and by $D^{*}$ its Hilbert adjoint. The associated Laplacian is the nonnegative self-adjoint operator
(2.9) $\Delta:=D^{*}D.$
(As $D$ is closed and densely defined, $D^{*}D$ is automatically self-adjoint and nonnegative.)
In the case where $\langle\cdot,\cdot\rangle$ is induced by a measure $\mu=\rho\,dx$ and $\widetilde{D}$$\nabla$, one recovers (up to sign conventions) the weighted Laplacian
$\Delta_{\mu}f=-\frac{1}{\rho}\,\mathrm{div}(\rho\nabla f),$
a classical object in the theory of Dirichlet forms and diffusion operators [5, 18].
2.4. Resolvents, norm-resolvent distance, and finite-degree compressions
Let $\Delta$ be a nonnegative self-adjoint operator on $H$. Its resolvent $(1+\Delta)^{-1}$ is a bounded operator on $H$ with $\|(1+\Delta)^{-1}\|\leq 1$. The norm-resolvent topology is defined by convergence in operator norm of resolvents,
$\Delta_{n}\to\Delta\quad\Longleftrightarrow\quad\|(1+\Delta_{n})^{-1}-(1+\Delta)^{-1}\|_{\mathcal{B}(H)}\to 0,$
see [8, 20].
For a fixed degree $N$, we will use two related finite-dimensional objects:
  • the truncated operator on $H_{\leq N}$,
    $\Delta^{(N)}:=P_{N}\,\Delta\big|_{H_{\leq N}}:H_{\leq N}\to H_{\leq N};$
  • the compressed resolvent on $H_{\leq N}$,
    $R^{\langle N\rangle}:=P_{N}(1+\Delta)^{-1}P_{N}\in\mathcal{B}(H_{\leq N}).$
In general $R^{\langle N\rangle}\neq(I+\Delta^{(N)})^{-1}$, unless $H_{\leq N}$ is invariant under $\Delta$. The stability results of the paper are formulated in terms of the resolvent and its compressions, which are the quantities directly controlled by norm-resolvent estimates.
2.5. Resolvents and operator topologies
Let $A$ be a self-adjoint, non-negative operator on a Hilbert space $H$. Its resolvent $(1+A)^{-1}$ is a bounded operator on $H$. A standard topology on the space of such operators is the norm-resolvent topology, defined by convergence in operator norm of resolvents:
$A_{n}\to A\quad\text{iff}\quad\|(1+A_{n})^{-1}-(1+A)^{-1}\|\to 0.$
This topology is metrizable and separated [8].
Remark 2.1.
When operators are associated with closed quadratic forms, norm-resolvent convergence is closely related to Mosco convergence of forms, a notion particularly suited to problems where the underlying measure or geometry varies [17, 11].
Remark 2.2 (Variable Hilbert spaces and quasi-unitary equivalence).
When comparing operators acting on different Hilbert spaces, one often introduces identification operators that are approximately unitary. This leads to the notion of quasi-unitary equivalence and generalized norm-resolvent convergence, developed notably in the context of spectral geometry and graph-like manifolds [19]. These tools allow one to compare spectral properties of operators defined on different spaces in a quantitative way.
All the notions recalled above are classical. In the following sections, they will be combined with the algebraic structure of polynomial spaces and the Gram–Schmidt orthogonalization procedure.
3. The Laplacian associated with a polynomial inner product
3.1. Degree growth and admissible derivations
Let $\mathcal{P}=\mathbb{K}[x_{1},\dots,x_{d}]$ be the algebra of real or complex polynomials, endowed with the standard filtration by total degree,
$\mathcal{P}_{\leq N}=\{p\in\mathcal{P}:\deg p\leq N\}.$
Any algebraic derivation
$D:\mathcal{P}\to\mathcal{P}$
admits a unique representation as a polynomial vector field,
(3.1) $D=\sum_{i=1}^{d}a_{i}(x)\,\partial_{x_{i}},\qquad a_{i}\in\mathcal{P}.$
A basic structural question concerns the interaction between the derivation $D$ and the degree filtration. Since each partial derivative $\partial_{x_{i}}$ lowers the degree by at most one, the behavior of $D$ with respect to the degree is governed by the polynomial coefficients $a_{i}$. More precisely, setting
$\delta(D):=\max_{1\leq i\leq d}\bigl(\deg a_{i}-1\bigr),$
one has, for every nonzero polynomial $p\in\mathcal{P}$,
(3.2) $\deg(Dp)\leq\deg(p)+\delta(D).$
In particular, the derivation $D$ may increase the degree of certain polynomials whenever $\delta(D)>0$, i.e. whenever at least one coefficient $a_{i}$ has degree greater than or equal to $2$. Conversely, $D$ satisfies
$\deg(Dp)\leq\deg(p)\quad\text{for all }p\in\mathcal{P}$
if and only if all coefficients $a_{i}$ are affine functions. In this case, $D$ generates an infinitesimal affine transformation of $\mathbb{R}^{d}$.
This observation has important consequences for the theory developed in this paper. Such derivations are incompatible with a graded orthogonalization procedure based on increasing degree and lead to Laplacian operators whose matrix representations are not locally finite with respect to the degree decomposition.
For this reason, throughout the present work we restrict attention to derivations which do not increase polynomial degree, and in particular to geometric derivations such as partial derivatives or gradients, for which $\delta(D)\leq 0$. In other words, we assume that $D$ restricts to a linear endomorphism of each vector space $\mathcal{P}_{\leq N}.$ This restriction ensures that the associated Laplacians interact in a controlled manner with the degree filtration
$\mathcal{P}_{\leq 1}\subset\mathcal{P}_{\leq 2}\subset\cdots\subset\mathcal{P}_{\leq N}\subset\mathcal{P}_{\leq N+1}\subset\cdots$
3.2. Closability of derivations and definition of $\Delta$
Let $\Omega\subset\mathbb{R}^{d}$ be a domain and let
$\mathcal{P}(\Omega):=\{\,p|_{\Omega}:\ p\in\mathbb{C}[x_{1},\dots,x_{d}]\,\}$
be the space of polynomial functions on $\Omega$. Let $\langle\cdot,\cdot\rangle$ be a positive semidefinite sesquilinear form on $\mathcal{P}(\Omega)$ (conjugate-linear in the first variable), and set
$N:=\{p\in\mathcal{P}(\Omega):\langle p,p\rangle=0\},\qquad\mathcal{H}_{0}:=\mathcal{P}(\Omega)/N,$
so that $\langle\cdot,\cdot\rangle$ induces an inner product on $\mathcal{H}_{0}$. Denote by $H$ the Hilbert completion of $(\mathcal{H}_{0},\langle\cdot,\cdot\rangle)$.
Derivations and descent to the quotient. Let $\widetilde{D}$ be an algebraic derivation on $\mathcal{P}(\Omega)$, typically $\widetilde{D}$$\nabla$ (or a finite family of derivations), taking values in $H^{m}$. We assume that
(3.3) $N\subset\ker\widetilde{D}$,
so that $\widetilde{D}$ induces a well-defined linear operator
$D_{0}:\mathcal{H}_{0}\to H^{m},\qquad D_{0}([p])=\widetilde{D}(p).$
Remark 3.1 (Positive definite case).
If $\langle\cdot,\cdot\rangle$ is positive definite on $\mathcal{P}(\Omega)$, then $N=\{0\}$ and the quotient step is unnecessary: one simply has $\mathcal{H}_{0}=\mathcal{P}(\Omega)$. In this situation (3.3) is automatic.
Closability and definition of the Laplacian. Assume that $D_{0}$ is densely defined in $H$ (which holds in all our main examples, since $\mathcal{H}_{0}$ is dense in $H$) and that $D_{0}$ is closable. We denote by
$D:=\overline{D_{0}}:\mathrm{Dom}(D)\subset H\to H^{m}$
its closure, which is a closed densely defined operator, and by $D^{*}$ its Hilbert adjoint. We then define the associated (positive) Laplacian by
(3.4) $\Delta:=D^{*}D.$
Since $D$ is closed and densely defined, the operator $D^{*}D$ is automatically self-adjoint and nonnegative.
Remarks on examples. In measure-induced settings (e.g. $\langle f,g\rangle=\int f\overline{g}\,d\mu$) and Sobolev-type settings, closability of $D_{0}$ is standard and can be formulated at the level of closed quadratic forms; see, e.g., [5, 18, 8, 17, 11].
3.3. Matrix representation in an orthonormal polynomial basis
Let $(P_{n})_{n\geq 0}$ be an orthonormal basis of polynomials in $H$, obtained by Gram–Schmidt orthogonalization of a canonical base $(E_{n})_{n\geq 0}$ of polynomials where indexation is non decreasing along the degree filtration.
Remark 3.2.
In several variables, the base $(E_{n})_{n\geq 0}$ is more naturally filtered with respect to a multi-index but the re-indexation with respect to $\mathbb{N}$ is aloways possible; therfore the discussion below is unchanged at the level of operator matrices.
Derivative matrix. Let us recall that we assume that $D:\mathrm{Dom}(D)\subset H$ restricts to a linear map $\mathcal{P}_{\leq N}\rightarrow\mathcal{P}_{\leq N}.$ Define the matrix coefficients of $D$ in the polynomial basis by
(3.5) $B_{k,n}:=\big\langle DP_{n},\,E_{k}\big\rangle.$
Since $(P_{n})$ is ordered by nondecreasing degree, the operator matrix of $D$ in polynomial coordinates is lower triangular ( or strictly lower triangular in the one-dimensional situation, with $D=\frac{d}{dx}$) or block-lower triangular (in a graded multi-index ordering).
Laplacian matrix. By definition $\Delta=D^{*}D$ is positive and self-adjoint, and its matrix in the orthonormal basis $(P_{n})$ is
(3.6) $\Delta_{m,n}:=\langle\Delta P_{n},P_{m}\rangle=\langle DP_{n},DP_{m}\rangle.$
If $D$ has matrix $B$ in polynomial coordinates, then
(3.7) $\Delta]=B^{*}B.$
In particular, $\Delta]$ is Hermitian and positive definite.
3.4. Banded recurrences and banded Laplacians
In many families of orthogonal polynomials, multiplication by coordinate functions has a banded representation in the orthonormal basis (three-term recurrence in one variable, block-banded in several variables). For Sobolev orthogonal polynomials, higher-order or banded recurrences are common, see the survey literature [16] and related algorithmic perspectives [2]. In the present setting, the key observation is:
Proposition 3.3 (Bandwidth propagation).
Let $(P_{n})_{n\geq 0}$ be an orthonormal polynomial basis of $H$ and let $B=(B_{k,n})_{k,n\geq 0}$ be the matrix of $D$ in this basis,
$DP_{n}=\sum_{k\geq 0}B_{k,n}\,P_{k},\qquad B_{k,n}:=\langle DP_{n},P_{k}\rangle.$
Assume that $B$ is banded with half-bandwidth $r\geq 1$, i.e.
$B_{k,n}=0\quad\text{whenever}\quad|k-n|>r.$
Then the matrix of $\Delta=D^{*}D$ in the same basis,
$\Delta_{m,n}:=\langle\Delta P_{n},P_{m}\rangle=\langle DP_{n},DP_{m}\rangle,$
is banded with half-bandwidth at most $2r$, i.e.
$\Delta_{m,n}=0\quad\text{whenever}\quad|m-n|>2r.$
Moreover $\Delta]$ is Hermitian and positive semidefinite.
Proof.
Since $(P_{n})$ is orthonormal, the coefficient array $B$ satisfies
$DP_{n}=\sum_{k\geq 0}B_{k,n}P_{k},\qquad\|DP_{n}\|^{2}=\sum_{k\geq 0}|B_{k,n}|^{2}.$
For $m,n\geq 0$ we compute, using orthonormality,
$\Delta_{m,n}=\langle DP_{n},DP_{m}\rangle$ $=\Big\langle\sum_{k\geq 0}B_{k,n}P_{k},\ \sum_{\ell\geq 0}B_{\ell,m}P_{\ell}\Big\rangle$
$=\sum_{k,\ell\geq 0}B_{k,n}\,\overline{B_{\ell,m}}\ \langle P_{k},P_{\ell}\rangle=\sum_{k\geq 0}B_{k,n}\,\overline{B_{k,m}}.$
This is exactly the matrix product identity
$\Delta]=B^{*}B,\qquad\Delta_{m,n}=\sum_{k\geq 0}\overline{B_{k,m}}\,B_{k,n}.$
Now assume |mn|>2r|m-n|>2r. If for some $k$ one has both $B_{k,n}\neq 0$ and $B_{k,m}\neq 0$, then by the bandedness hypothesis we must have $|k-n|\leq r$ and $|k-m|\leq r$, hence
$|m-n|\leq|m-k|+|k-n|\leq r+r=2r,$
a contradiction. Therefore for every $k$ at least one factor in $B_{k,n}\,\overline{B_{k,m}}$ is zero, and the sum vanishes: $\Delta_{m,n}=0$. This proves the half-bandwidth bound $2r$.
Finally, $\Delta]=B^{*}B$ implies that $\Delta]$ is Hermitian and positive semidefinite, since for any finitely supported vector $c=(c_{n})$,
$\sum_{m,n}\Delta_{m,n}c_{n}\overline{c_{m}}=\langle Bc,Bc\rangle_{\ell^{2}}\geq 0.$
Remark 3.4.
The hypothesis “$D$ banded” can be verified in concrete situations either by explicit connection relations (derivatives expanded in the orthonormal basis) or via structural results (semiclassical weights, coherent pairs, etc.) in the Sobolev orthogonality literature; see [16].
3.5. Example: thin annuli and mode-by-mode block structure
We now record the operator-structural consequences in the thin annulus setting treated in [13], where the orthogonal polynomials are constructed on a planar annulus with a (possibly fractional) Sobolev inner product and the analysis proceeds mode by mode with respect to the angular Fourier decomposition.
Let $\Omega_{\varepsilon}=\{(r,\theta):r\in(r_{0},r_{0}+\varepsilon),\ \theta\in[0,2\pi)\}$ be a thin annulus. In polar coordinates, the natural first-order derivation is $D=\nabla$. The rotational symmetry implies that the polynomial (pre-)Hilbert space decomposes orthogonally as a direct sum of Fourier modes (angular frequencies). In particular, the Gram–Schmidt orthogonal basis can be chosen so that each basis element carries a definite angular frequency (a standard separation principle on rotationally invariant domains, used explicitly in [13]).
Proposition 3.5 (Block diagonalization by angular modes).
In the setting of [13], the Hilbert space completion admits an orthogonal decomposition
$H\;=\;\bigoplus_{k\in\mathbb{Z}}H^{(k)},$
where $H^{(k)}$ is the closed subspace generated by polynomials with angular dependence $e^{ik\theta}$. Moreover, with respect to an orthonormal basis adapted to this decomposition, the Laplacian $\Delta=D^{*}D$ is block diagonal:
$\Delta\;=\;\bigoplus_{k\in\mathbb{Z}}\Delta^{(k)}.$
The proof can be decomposed into successive lemmas.
Lemma 3.6 (Angular decomposition induced by rotational invariance).
Let $A_{\varepsilon}=\{x\in\mathbb{R}^{2}:1-\varepsilon<|x|^{2}<1+\varepsilon\}$ and let $\langle\cdot,\cdot\rangle_{\varepsilon}$ be a (pre-)Hilbert inner product on $\mathcal{P}(A_{\varepsilon})$ whose completion is a Hilbert space $H_{\varepsilon}$. Assume that $\langle\cdot,\cdot\rangle_{\varepsilon}$ is rotationally invariant in the sense that for every rotation $R_{\phi}\in SO(2)$,
(3.8) $\langle f\circ R_{\phi},g\circ R_{\phi}\rangle_{\varepsilon}=\langle f,g\rangle_{\varepsilon},\qquad\forall f,g\in\mathcal{P}(A_{\varepsilon}),\ \forall\phi\in\mathbb{R}.$
Then the operators $U_{\phi}:H_{\varepsilon}\to H_{\varepsilon}$ defined by $U_{\phi}f:=f\circ R_{\phi}$ extend to a strongly continuous unitary representation of $S^{1}$, and $H_{\varepsilon}$ admits the orthogonal decomposition
$H_{\varepsilon}=\widehat{\bigoplus_{k\in\mathbb{Z}}}\,H_{\varepsilon}^{(k)},\qquad H_{\varepsilon}^{(k)}:=\{f\in H_{\varepsilon}:\ U_{\phi}f=e^{ik\phi}f\ \text{for all }\phi\}.$
Moreover, the orthogonal projection onto $H_{\varepsilon}^{(k)}$ is given by the Fourier projector
(3.9) $P_{k}f:=\frac{1}{2\pi}\int_{0}^{2\pi}e^{-ik\phi}\,U_{\phi}f\,d\phi,\qquad f\in H_{\varepsilon},$
where the integral is understood in the Bochner sense in $H_{\varepsilon}$.
Proof.
Step 1: unitary representation. By (3.8), for all polynomials $f,g$,
$\langle U_{\phi}f,U_{\phi}g\rangle_{\varepsilon}=\langle f,g\rangle_{\varepsilon},$
hence each $U_{\phi}$ extends by density to an isometry on $H_{\varepsilon}$. Since $U_{\phi}^{-1}=U_{-\phi}$, it is unitary. The group property $U_{\phi}U_{\psi}=U_{\phi+\psi}$ holds on $\mathcal{P}(A_{\varepsilon})$ and extends to $H_{\varepsilon}$.
Step 2: strong continuity. In the concrete Sobolev or fractional Sobolev settings used in [13], the map $\phi\mapsto U_{\phi}f$ is continuous in the corresponding norm for $f\in\mathcal{P}(A_{\varepsilon})$, because rotations act continuously on $L^{2}$ and on the relevant Sobolev seminorms (derivatives or fractional energies are rotation invariant). By density, the representation is strongly continuous on $H_{\varepsilon}$.
Step 3: Fourier projectors. Define $P_{k}$ by (3.9). Since $U_{\phi}$ is unitary and strongly continuous, the Bochner integral exists and $P_{k}$ is bounded with $\|P_{k}\|\leq 1$. A direct computation using $U_{\phi}U_{\psi}=U_{\phi+\psi}$ yields, for all $k,\ell\in\mathbb{Z}$,
$P_{k}P_{\ell}=\delta_{k\ell}P_{k},$
so $P_{k}$ is an orthogonal projection and the ranges are pairwise orthogonal.
Step 4: decomposition. For each $f\in H_{\varepsilon}$, the family of partial Fourier sums $S_{N}f:=\sum_{|k|\leq N}P_{k}f$ converges to $f$ in $H_{\varepsilon}$. This is the standard $L^{2}$-Fourier convergence argument applied to the unitary representation of the compact abelian group $S^{1}$ (here implemented concretely by (3.9)). Thus $H_{\varepsilon}$ is the Hilbert direct sum of the closed subspaces $H_{\varepsilon}^{(k)}:=Ran(P_{k})$. Finally, one checks that $Im(P_{k})$ coincides with the $k$-eigenspace of the representation: indeed, for $f\in H_{\varepsilon}$,
$U_{\psi}(P_{k}f)=\frac{1}{2\pi}\int_{0}^{2\pi}e^{-ik\phi}\,U_{\psi+\phi}f\,d\phi=e^{ik\psi}\frac{1}{2\pi}\int_{0}^{2\pi}e^{-ik\phi^{\prime}}\,U_{\phi^{\prime}}f\,d\phi^{\prime}=e^{ik\psi}P_{k}f,$
so $Im(P_{k})\subset H_{\varepsilon}^{(k)}$, and conversely if $U_{\psi}f=e^{ik\psi}f$ then $P_{k}f=f$ and $P_{\ell}f=0$ for $\ell\neq k$. This proves the claim. ∎
Lemma 3.7 (Intertwining of $\nabla$ with rotations and block diagonalization of $\Delta$).
Assume the hypotheses of Lemma 3.6 and let $D=\nabla$. Assume in addition that the inner product $\langle\cdot,\cdot\rangle_{\varepsilon}$ is of Sobolev type as in [13], so that $D$ extends to a closable operator $D:Dom(D)\subset H_{\varepsilon}\to H_{\varepsilon}^{\mathrm{vec}}$, where $H_{\varepsilon}^{\mathrm{vec}}$ is a Hilbert space of vector fields on $A_{\varepsilon}$ (for instance $L^{2}$- or Sobolev-type), and that the vector-field inner product is also rotation invariant.
Let $U_{\phi}^{\mathrm{vec}}$ act on vector fields by
$(U_{\phi}^{\mathrm{vec}}F)(x):=R_{\phi}\,F(R_{\phi}x),$
which is unitary on $H_{\varepsilon}^{\mathrm{vec}}$ by rotational invariance. Then for all $\phi$ and all $f\in Dom(D)$,
(3.10) $D(U_{\phi}f)=U_{\phi}^{\mathrm{vec}}(Df).$
Consequently, the Laplacian $\Delta:=D^{*}D$ commutes with $U_{\phi}$,
(3.11) $U_{\phi}\Delta=\Delta U_{\phi},$
and each angular subspace $H_{\varepsilon}^{(k)}$ is invariant under $\Delta$. In particular, $\Delta$ is block diagonal with respect to the decomposition $H_{\varepsilon}=\widehat{\bigoplus}_{k\in\mathbb{Z}}H_{\varepsilon}^{(k)}$.
Proof.
Step 1: chain rule (intertwining). For smooth functions (hence for polynomials) the chain rule gives
$\nabla(f\circ R_{\phi})(x)=(DR_{\phi})^{\!\top}\,(\nabla f)(R_{\phi}x).$
Since $R_{\phi}$ is an orthogonal matrix, $(DR_{\phi})^{\!\top}=R_{\phi}^{\top}=R_{-\phi}$. Equivalently,
$\nabla(f\circ R_{\phi})(x)=R_{-\phi}\,(\nabla f)(R_{\phi}x).$
Rewriting this in terms of $U_{\phi}^{\mathrm{vec}}$ as defined above gives exactly (3.10) on $\mathcal{P}(A_{\varepsilon})$, hence on $Dom(D)$ by density and closability.
Step 2: commutation of $\Delta$ with rotations. Let $f\in Dom(\Delta)$ and $g\in Dom(D)$. Using unitarity of $U_{\phi}$ and $U_{\phi}^{\mathrm{vec}}$, together with (3.10), we compute
$\langle D(U_{\phi}f),Dg\rangle_{\mathrm{vec}}=\langle U_{\phi}^{\mathrm{vec}}Df,Dg\rangle_{\mathrm{vec}}=\langle Df,(U_{\phi}^{\mathrm{vec}})^{-1}Dg\rangle_{\mathrm{vec}}=\langle Df,D(U_{-\phi}g)\rangle_{\mathrm{vec}}=\langle\Delta f,U_{-\phi}g\rangle=\langle U_{\phi}\Delta f,g\rangle.$
By definition of the adjoint and density of $Dom(D)$, this shows that $\Delta(U_{\phi}f)=U_{\phi}\Delta f$, i.e. (3.11).
Step 3: invariance of modes. If $f\in H_{\varepsilon}^{(k)}$, then $U_{\phi}f=e^{ik\phi}f$ for all $\phi$. Applying (3.11) yields
$U_{\phi}(\Delta f)=\Delta(U_{\phi}f)=\Delta(e^{ik\phi}f)=e^{ik\phi}\Delta f,$
hence $\Delta f\in H_{\varepsilon}^{(k)}$. Therefore each $H_{\varepsilon}^{(k)}$ is invariant and $\Delta$ is block diagonal. ∎
Remark 3.8 (Bandedness within each mode).
A central output of [13] is the existence of banded recurrences for the Sobolev orthogonal polynomials obtained in each mode, together with asymptotics as $\varepsilon\to 0$. When the derivative/connection relations within a fixed mode yield a banded representation of $D$ in the mode-adapted orthonormal basis, Proposition 3.3 implies that each block matrix $\Delta^{(k)}]$ is itself banded (with an explicit bandwidth bound in terms of the recurrence width). This provides a concrete and numerically tractable matrix model for the spectral study of $\Delta$ in the thin-annulus Sobolev setting.
4. The Laplacian in the classical one-dimensional weighted $L^{2}$ setting
4.1. Weighted $L^{2}$ space and adjoint of the derivative
Let $I\subset\mathbb{R}$ be a (bounded or unbounded) interval and let $\mu$ be a finite Radon measure on $I$ which is absolutely continuous with respect to Lebesgue measure:
$d\mu(x)=w(x)\,dx,\qquad w\geq 0,\qquad\int_{I}w(x)\,dx<\infty.$
We consider the Hilbert space $H:=L^{2}(I,w\,dx)$ with inner product
$\langle f,g\rangle_{w}:=\int_{I}f(x)\,\overline{g(x)}\,w(x)\,dx.$
Assume that all moments exist (at least up to the degrees considered) so that the Gram–Schmidt procedure applied to $(1,x,x^{2},\dots)$ produces an orthonormal polynomial family $(p_{n})_{n\geq 0}$ in $H$ (see, e.g., [23, 3]).
Let $D=\partial_{x}$ on $C_{c}^{\infty}(I)$ (or on polynomials, viewed as a dense subspace of $H$ whenever it makes sense). A formal computation yields the weighted adjoint:
(4.1) $\langle f^{\prime},g\rangle_{w}=-\langle f,g^{\prime}+(w^{\prime}/w)g\rangle_{w}+\big[f\,\overline{g}\,w\big]_{\partial I},$
whenever the boundary term is meaningful. Thus, on a domain enforcing $\big[f\,\overline{g}\,w\big]_{\partial I}=0$ (e.g. compact support, or suitable weighted boundary conditions), the adjoint of $D$ is
$D^{*}g=-g^{\prime}-\frac{w^{\prime}}{w}\,g.$
Consequently, the associated positive Laplacian is the (weighted) Sturm–Liouville operator
(4.2) $\Delta_{w}:=D^{*}D=-\frac{1}{w}\frac{d}{dx}\Big(w\,\frac{d}{dx}\Big)=-\frac{d^{2}}{dx^{2}}-\frac{w^{\prime}}{w}\,\frac{d}{dx}.$
Under standard hypotheses (closability of $D$, choice of Friedrichs extension, etc.) $\Delta_{w}$ is a nonnegative self-adjoint operator on $H$; this is routine in the operator/form frameworks (see [8, 18]).
4.2. Matrix representation in the orthonormal polynomial basis
Let $(p_{n})_{n\geq 0}$ be the orthonormal polynomials in $L^{2}(I,w\,dx)$. The Laplacian matrix in this basis is given by
(4.3) $(\Delta_{w})_{mn}:=\langle\Delta_{w}p_{n},p_{m}\rangle_{w}=\langle p_{n}^{\prime},p_{m}^{\prime}\rangle_{w}=\int_{I}p_{n}^{\prime}(x)\,\overline{p_{m}^{\prime}(x)}\,w(x)\,dx,$
where the last equality uses the definition $\Delta_{w}=D^{*}D$ and the Hilbert structure.
Equivalently, if one expands derivatives in the polynomial basis,
(4.4) $p_{n}^{\prime}(x)=\sum_{k=0}^{n-1}b_{n,k}\,p_{k}(x),\qquad b_{n,k}:=\langle p_{n}^{\prime},p_{k}\rangle_{w},$
then the Laplacian matrix is the Gram matrix of the derivative map:
(4.5) $(\Delta_{w})_{mn}=\sum_{k\geq 0}b_{n,k}\,\overline{b_{m,k}},\qquad\text{i.e.}\qquad[\Delta_{w}]=B^{*}B,$
where $B=(b_{n,k})$ is (strictly) lower triangular by degree.
4.3. Sparsity and classical weights
In general, the derivative expansion (4.4) is not banded: $p_{n}^{\prime}$ may involve many lower modes. However, for the classical families (Jacobi, Laguerre, Hermite), one has strong “lowering” relations and second-order differential equations of the form
(4.6) $\sigma(x)\,p_{n}^{\prime\prime}(x)+\tau(x)\,p_{n}^{\prime}(x)=\lambda_{n}\,p_{n}(x),$
with $\deg\sigma\leq 2$ and $\deg\tau\leq 1$ (see [23, 3]). In those cases, the operator in (4.2) (or a closely related conjugate/operator with polynomial coefficients) acts diagonally on the polynomial basis, so the Laplacian matrix is (diagonal or at least finite-band after a fixed change of basis). This is the one-dimensional analogue of the finite-band phenomena encountered in Sobolev inner products.
4.4. Relation with the Jacobi matrix and commutator structure
Let $(p_{n})_{n\geq 0}$ be the orthonormal polynomial basis in $L^{2}(I,w\,dx)$. Multiplication by the coordinate function defines a symmetric operator
$(M_{x}f)(x)=xf(x),$
whose matrix in the basis $(p_{n})$ is the Jacobi matrix
(4.7) $x\,p_{n}(x)=a_{n+1}p_{n+1}(x)+b_{n}p_{n}(x)+a_{n}p_{n-1}(x),\qquad a_{n}>0,\;b_{n}\in\mathbb{R}.$
The Jacobi matrix $J$ is tridiagonal and self-adjoint on $\ell^{2}(\mathbb{N})$, and its spectral measure coincides with $\mu=w(x)\,dx$ (Favard’s theorem; see [23, 3, 21]).
Derivative–multiplication commutator. On smooth functions one has the fundamental commutator identity
(4.8) $D,M_{x}]=I,$
where $D=\partial_{x}$. Passing to adjoints with respect to the weighted inner product yields
$D^{*},M_{x}]=-I-\frac{w^{\prime}}{w}M_{x}+M_{x}\frac{w^{\prime}}{w},$
so that the weighted geometry introduces lower-order correction terms unless $w$ is constant.
Laplacian expressed via Jacobi data. Using $\Delta_{w}=D^{*}D$, one computes the commutator
(4.9) $\Delta_{w},M_{x}]=D^{*}[D,M_{x}]+[D^{*},M_{x}]D=D^{*}+[D^{*},M_{x}]D,$
which is a first-order differential operator. In the orthonormal polynomial basis, this identity implies that the matrix of $\Delta_{w}$ is controlled by the Jacobi coefficients $(a_{n},b_{n})$ together with the expansion coefficients of $p_{n}^{\prime}$ in the basis $(p_{k})$.
More explicitly, since
$p_{n}^{\prime}=\sum_{k=0}^{n-1}b_{n,k}\,p_{k},$
one has
$(\Delta_{w})_{mn}=\sum_{k=0}^{\min(m,n)-1}b_{n,k}\,\overline{b_{m,k}},$
while the Jacobi matrix governs multiplication by $x$ through (4.7). Thus, the pair $(J,\Delta_{w})$ encodes simultaneously:
  • the three-term recurrence (algebraic structure);
  • the energy form $\int|f^{\prime}|^{2}w\,dx$ (geometric structure).
Classical weights. For the classical weights (Jacobi, Laguerre, Hermite), the orthonormal polynomials are eigenfunctions of a second-order differential operator
$\mathcal{L}=\sigma(x)\partial_{x}^{2}+\tau(x)\partial_{x},\qquad\deg\sigma\leq 2,\ \deg\tau\leq 1,$
which is unitarily equivalent to $\Delta_{w}$ up to a multiplication operator. In this situation, the polynomial basis diagonalizes $\mathcal{L}$, and hence the Laplacian matrix is diagonal (or block-diagonal after a fixed normalization), reflecting the complete integrability of the classical case [23, 3].
Perspective. Outside the classical setting, $\Delta_{w}$ and $J$ no longer commute, and the Laplacian matrix becomes genuinely non-diagonal. The deviation from diagonality provides a quantitative measure of how far the orthogonalization procedure is from the classical Sturm–Liouville regime. This observation will be central in comparing different polynomial geometries via their associated Laplacians.
5. Polynomial Hilbert geometries and associated Laplacians
5.1. Polynomial Hilbert geometries
Let $\Omega\subset\mathbb{R}^{d}$ be a domain and let $\mathcal{P}(\Omega)$ denote the space of polynomial functions on $\Omega$ (i.e. restrictions of ambient polynomials). Building on the explicitly treated examples, we can now introduce a unifying geometric framework for polynomial inner products and their associated operators. Some overlap with the preliminary section is therefore intentional: we prefer to restate the key constructions in a self-contained manner, in order to keep the general setting clear and easy to follow.
Definition 5.1 (Polynomial Hilbert geometry).
Let $\Omega\subset\mathbb{R}^{d}$ and let $\mathcal{P}(\Omega)$ be the space of (restrictions to $\Omega$ of) complex polynomials. A polynomial Hilbert geometry on $\Omega$ is a triple
$\mathfrak{G}=(\mathcal{P}(\Omega),\langle\cdot,\cdot\rangle,D),$
where:
  • $\langle\cdot,\cdot\rangle$ is a positive semidefinite sesquilinear form on $\mathcal{P}(\Omega)$ (conjugate-linear in the first variable);
  • $N:=\{p\in\mathcal{P}(\Omega):\langle p,p\rangle=0\}$ and $\mathcal{H}_{0}:=\mathcal{P}(\Omega)/N$; the form $\langle\cdot,\cdot\rangle$ descends to a Hermitian inner product on $\mathcal{H}_{0}$;
  • $H$ denotes the Hilbert completion of $(\mathcal{H}_{0},\langle\cdot,\cdot\rangle)$;
  • $\widetilde{D}:\mathcal{P}(\Omega)\to H^{m}$ is a derivation (typically $\widetilde{D}$$\nabla$), and we assume $N\subset\ker\widetilde{D}$ so that $\widetilde{D}$ induces a well-defined linear map $D_{0}:\mathcal{H}_{0}\to H^{m}$ by $D_{0}([p])=\widetilde{D}(p)$. We assume $D_{0}$ is densely defined and closable.
We denote by $D:=\overline{D_{0}}$ its closure and set $\Delta:=D^{*}D$.
Throughout, we denote again by $D$ its closure and by $D^{*}$ its Hilbert adjoint.
Remark 5.2 (On the notion of domain and the role of the quotient).
In this paper, a “domain” is simply a subset $\Omega\subset\mathbb{R}^{d}$, not necessarily open. We always interpret
$\mathcal{P}(\Omega):=\{\,p|_{\Omega}:p\in\mathbb{C}[x_{1},\dots,x_{d}]\,\}$
as the space of polynomial functions on $\Omega$ (restrictions of ambient polynomials). When $\Omega$ has empty interior (or more generally when the restriction map $p\mapsto p|_{\Omega}$ is not injective), distinct polynomials may induce the same function on $\Omega$. For instance, for $\Omega=S^{1}\subset\mathbb{R}^{2}$ one has the relation $x^{2}+y^{2}=1$ on $\Omega$, so the ideal $\langle x^{2}+y^{2}-1\rangle$ lies in the kernel of any $L^{2}(S^{1})$-type inner product. This explains the need to pass to the quotient $\mathcal{P}(\Omega)/\ker\langle\cdot,\cdot\rangle$.
Definition 5.3 (Associated Laplacian).
The Laplacian associated with $\mathfrak{G}$ is the positive self-adjoint operator
$\Delta_{\mathfrak{G}}:=D^{*}D$
defined on $H$.
Different polynomial Hilbert geometries may induce the same operator-theoretic structure.
Definition 5.4 (Unitary equivalence).
Two polynomial Hilbert geometries $\mathfrak{G}=(\mathcal{P}(\Omega),\langle\cdot,\cdot\rangle,D)$ and $\mathfrak{G}^{\prime}=(\mathcal{P}(\Omega),\langle\cdot,\cdot\rangle^{\prime},D^{\prime})$ are said to be unitarily equivalent if there exists a unitary operator $U:H\to H^{\prime}$ such that:
  1. (1)
    $U(\mathcal{H}_{0})=\mathcal{H}_{0}^{\prime}$ (identified as dense subspaces of $H$ and $H^{\prime}$);
  2. (2)
    $UD=D^{\prime}U$ on $\mathcal{H}_{0}$.
Unitary equivalence implies
$U\Delta_{\mathfrak{G}}U^{-1}=\Delta_{\mathfrak{G}^{\prime}}.$
5.2. Resolvent-based distance
We now introduce the operator-theoretic distance which will be used throughout the paper.
Definition 5.5 (Resolvent distance).
Let $\mathfrak{G}_{1},\mathfrak{G}_{2}$ be two polynomial Hilbert geometries whose Laplacians $\Delta_{1},\Delta_{2}$ act on the same Hilbert space $H$ (or are identified via a fixed unitary equivalence). We define
(5.1) $d_{\mathrm{res}}(\mathfrak{G}_{1},\mathfrak{G}_{2}):=\big\|(1+\Delta_{1})^{-1}-(1+\Delta_{2})^{-1}\big\|_{\mathcal{B}(H)}.$
This is the classical norm-resolvent distance between nonnegative self-adjoint operators [8, 18].
Proposition 5.6.
Let $H$ be a Hilbert space and let $\Delta_{1},\Delta_{2}$ be nonnegative self-adjoint operators on $H$. Define
$d_{\mathrm{res}}(\Delta_{1},\Delta_{2}):=\big\|(1+\Delta_{1})^{-1}-(1+\Delta_{2})^{-1}\big\|_{\mathcal{B}(H)}.$
Then $d_{\mathrm{res}}$ is a metric on the set of nonnegative self-adjoint operators on $H$. Moreover, for every unitary $U\in\mathcal{U}(H)$,
$d_{\mathrm{res}}(U\Delta_{1}U^{-1},\,U\Delta_{2}U^{-1})=d_{\mathrm{res}}(\Delta_{1},\Delta_{2}).$
Proof.
Step 0: boundedness of resolvents. Since each $\Delta_{i}$ is self-adjoint and nonnegative, its spectrum satisfies $\sigma(\Delta_{i})\subset[0,\infty)$. Hence $1\notin\sigma(\Delta_{i})$ and the resolvent
$R_{i}:=(1+\Delta_{i})^{-1}$
exists as a bounded operator on $H$. Moreover, by the spectral theorem,
$\|R_{i}\|_{\mathcal{B}(H)}=\sup_{\lambda\in\sigma(\Delta_{i})}\frac{1}{1+\lambda}\leq 1.$
Therefore $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=\|R_{1}-R_{2}\|$ is well-defined and finite.
Step 1: positivity and symmetry. By definition, $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})$ is the norm of a bounded operator, hence $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})\geq 0$. Also,
$d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=\|R_{1}-R_{2}\|=\|-(R_{2}-R_{1})\|=d_{\mathrm{res}}(\Delta_{2},\Delta_{1}),$
so $d_{\mathrm{res}}$ is symmetric.
Step 2: triangle inequality. Let $\Delta_{1},\Delta_{2},\Delta_{3}$ be nonnegative self-adjoint on $H$ and set $R_{i}=(1+\Delta_{i})^{-1}$. Then
$R_{1}-R_{3}=(R_{1}-R_{2})+(R_{2}-R_{3}),$
and by the triangle inequality in the operator norm,
$d_{\mathrm{res}}(\Delta_{1},\Delta_{3})=\|R_{1}-R_{3}\|\leq\|R_{1}-R_{2}\|+\|R_{2}-R_{3}\|=d_{\mathrm{res}}(\Delta_{1},\Delta_{2})+d_{\mathrm{res}}(\Delta_{2},\Delta_{3}).$
Step 3: separation (identity of indiscernibles). Assume $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=0$. Then $\|R_{1}-R_{2}\|=0$, hence $R_{1}=R_{2}$ as bounded operators. We claim that this implies $\Delta_{1}=\Delta_{2}$.
Indeed, since $\Delta_{i}$ is self-adjoint, the operator $1+\Delta_{i}$ is self-adjoint with domain $\mathrm{Dom}(\Delta_{i})$, and it is bijective from $\mathrm{Dom}(\Delta_{i})$ onto $H$, with bounded inverse $R_{i}$. In particular,
$(1+\Delta_{i})R_{i}=I\quad\text{on }H,\qquad R_{i}(1+\Delta_{i})=I\quad\text{on }\mathrm{Dom}(\Delta_{i}).$
Since $R_{1}=R_{2}=:R$, we have $\mathrm{Ran}(R)=\mathrm{Dom}(\Delta_{1})=\mathrm{Dom}(\Delta_{2})$. Moreover, on this common domain,
$(1+\Delta_{1})u=R^{-1}u=(1+\Delta_{2})u.$
Hence $\Delta_{1}u=\Delta_{2}u$ for all $u\in\mathrm{Dom}(\Delta_{1})=\mathrm{Dom}(\Delta_{2})$, so $\Delta_{1}=\Delta_{2}$. Therefore $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=0$ implies $\Delta_{1}=\Delta_{2}$.
Conversely, if $\Delta_{1}=\Delta_{2}$, then $R_{1}=R_{2}$ and $d_{\mathrm{res}}(\Delta_{1},\Delta_{2})=0$. This proves separation.
Step 4: unitary invariance. Let $U$ be unitary on $H$. If $\Delta$ is self-adjoint and nonnegative, then so is $\widetilde{\Delta}:=U\Delta U^{-1}$, with $\sigma(\widetilde{\Delta})=\sigma(\Delta)$. Moreover,
$1+\widetilde{\Delta}=U(1+\Delta)U^{-1},$
hence by taking inverses on both sides,
$(1+\widetilde{\Delta})^{-1}=U(1+\Delta)^{-1}U^{-1}.$
Applying this to $\Delta_{1},\Delta_{2}$ gives
$d_{\mathrm{res}}(U\Delta_{1}U^{-1},\,U\Delta_{2}U^{-1})$ $=\big\|U(1+\Delta_{1})^{-1}U^{-1}-U(1+\Delta_{2})^{-1}U^{-1}\big\|$
$=\big\|U\big((1+\Delta_{1})^{-1}-(1+\Delta_{2})^{-1}\big)U^{-1}\big\|.$
Since conjugation by a unitary preserves the operator norm, $\|UTU^{-1}\|=\|T\|$, we obtain
$d_{\mathrm{res}}(U\Delta_{1}U^{-1},\,U\Delta_{2}U^{-1})=d_{\mathrm{res}}(\Delta_{1},\Delta_{2}).$
This completes the proof that $d_{\mathrm{res}}$ is a metric and is unitarily invariant. ∎
Proof.
If $d_{\mathrm{res}}(\mathfrak{G}_{1},\mathfrak{G}_{2})=0$, then $(1+\Delta_{1})^{-1}=(1+\Delta_{2})^{-1}$, hence $\Delta_{1}=\Delta_{2}$ by functional calculus. Unitary equivalence preserves resolvents, and the norm induces a metric structure. ∎
Remark 5.7.
When comparing two polynomial Hilbert geometries, we will always assume that their Laplacians are realized on the same Hilbert space (or transported to a common Hilbert space through a fixed unitary identification). In that case $d_{\mathrm{res}}$ is the usual norm-resolvent distance between the transported Laplacians. We do not claim that it separates geometries inducing the same Laplacian.
5.3. Finite-degree truncations
Let $\mathcal{P}_{\leq N}(\Omega)$ denote the space of polynomials of degree at most $N$, and let $H_{\leq N}\subset H$ be its (finite-dimensional) image in $H$. We denote by $P_{N}:H\to H_{\leq N}$ the orthogonal projector.
Definition 5.8 (Truncated Laplacian).
The truncated Laplacian of order $N$ associated with $\mathfrak{G}$ is the operator on $H_{\leq N}$
$\Delta_{\mathfrak{G}}^{(N)}:=P_{N}\,\Delta_{\mathfrak{G}}\big|_{H_{\leq N}}:H_{\leq N}\to H_{\leq N}.$
Definition 5.9 (Compressed resolvent).
We also consider the compressed resolvent on $H_{\leq N}$:
$R_{\mathfrak{G}}^{\langle N\rangle}:=P_{N}(1+\Delta_{\mathfrak{G}})^{-1}P_{N}\in\mathcal{B}(H_{\leq N}).$
Remark 5.10.
In general one has $R_{\mathfrak{G}}^{\langle N\rangle}\neq(I+\Delta_{\mathfrak{G}}^{(N)})^{-1}$. They coincide when $H_{\leq N}$ is invariant under $\Delta_{\mathfrak{G}}$ (e.g. if the derivation preserves the degree filtration).
Theorem 5.11 (Finite-degree stability).
Let $\mathfrak{G}_{1},\mathfrak{G}_{2}$ be two polynomial Hilbert geometries, realized on a common Hilbert space $H$ via a fixed unitary identification, and let
$R_{i}:=(1+\Delta_{\mathfrak{G}_{i}})^{-1}\in\mathcal{B}(H),\qquad i=1,2.$
Assume that
$d_{\mathrm{res}}(\mathfrak{G}_{1},\mathfrak{G}_{2})=\|R_{1}-R_{2}\|<\varepsilon.$
Fix $N\in\mathbb{N}$ and let $P_{N}$ denote the orthogonal projection onto the closed subspace $H_{\leq N}\subset H$ generated by polynomials of degree $\leq N$. Define the truncated resolvents by
$R_{i}^{(N)}:=P_{N}R_{i}P_{N}\in\mathcal{B}(H_{\leq N}).$
Then
(5.2) $\|R_{1}^{(N)}-R_{2}^{(N)}\|\leq\varepsilon,$
so the conclusion holds with $C_{N}=1$.
Moreover, since $H_{\leq N}$ is finite-dimensional, the eigenvalues of $R_{1}^{(N)}$ and $R_{2}^{(N)}$ (and hence the spectral data of the compressed operators $\Delta_{\mathfrak{G}_{i}}^{(N)}:=P_{N}\Delta_{\mathfrak{G}_{i}}P_{N}$) and the matrix coefficients of $R_{i}^{(N)}$ in any fixed orthonormal basis of $H_{\leq N}$ are Lipschitz close, with constants depending only on $\dim(H_{\leq N})$.
Proof.
Step 1: resolvent control on the truncated space. Since $P_{N}$ is an orthogonal projection, $\|P_{N}\|=1$ and $P_{N}^{2}=P_{N}$. Therefore,
$R_{1}^{(N)}-R_{2}^{(N)}=P_{N}(R_{1}-R_{2})P_{N},$
and taking operator norms yields
$\|R_{1}^{(N)}-R_{2}^{(N)}\|\leq\|P_{N}\|\,\|R_{1}-R_{2}\|\,\|P_{N}\|=\|R_{1}-R_{2}\|<\varepsilon.$
This proves (5.2).
Step 2: Lipschitz control of spectral data. The space $H_{\leq N}$ has finite dimension $m=\dim(H_{\leq N})$. Both $R_{1}^{(N)}$ and $R_{2}^{(N)}$ are bounded self-adjoint operators on $H_{\leq N}$. Let $\lambda_{1}^{(i)}\geq\cdots\geq\lambda_{m}^{(i)}$ be the eigenvalues of $R_{i}^{(N)}$ (counted with multiplicity). By Weyl’s perturbation inequality for Hermitian matrices (finite-dimensional spectral perturbation theory),
$\max_{1\leq j\leq m}|\lambda_{j}^{(1)}-\lambda_{j}^{(2)}|\leq\|R_{1}^{(N)}-R_{2}^{(N)}\|\leq\varepsilon.$
In particular the spectra of the truncated resolvents are Lipschitz close.
Since $\Delta_{\mathfrak{G}_{i}}^{(N)}=P_{N}\Delta_{\mathfrak{G}_{i}}P_{N}$ is self-adjoint and nonnegative on $H_{\leq N}$, its eigenvalues $\mu_{j}^{(i)}\geq 0$ are related to those of $R_{i}^{(N)}$ by
$\lambda_{j}^{(i)}=\frac{1}{1+\mu_{j}^{(i)}}.$
Hence, whenever one has an a priori upper bound $\mu_{j}^{(i)}\leq M_{N}$ (which holds automatically on the fixed finite-dimensional space $H_{\leq N}$), the map $t\mapsto(1+t)^{-1}$ is Lipschitz on $0,M_{N}]$ with constant $\leq 1$, and the eigenvalues $\mu_{j}^{(i)}$ are Lipschitz close as well, with a constant depending only on $M_{N}$ (and thus only on $N$ and the two geometries restricted to $H_{\leq N}$).
Step 3: Lipschitz control of matrix coefficients. Fix any orthonormal basis $(e_{1},\dots,e_{m})$ of $H_{\leq N}$. Then the matrix coefficients satisfy
$|\langle(R_{1}^{(N)}-R_{2}^{(N)})e_{\alpha},e_{\beta}\rangle|\leq\|R_{1}^{(N)}-R_{2}^{(N)}\|\leq\varepsilon,\qquad 1\leq\alpha,\beta\leq m,$
so all coefficients are Lipschitz close.
This proves the theorem. ∎
6. Stability of polynomial projections and Gram–Schmidt orthogonalization
Let $\mathfrak{G}=(\mathcal{P}(\Omega),\langle\cdot,\cdot\rangle,D)$ be a polynomial Hilbert geometry, with associated Hilbert space $H$ and Laplacian $\Delta$. For $N\geq 0$, the projector $P_{N}$ depends on the geometry $\mathfrak{G}$ through the inner product. We write $P_{N}^{(i)}$ when several geometries $\mathfrak{G}_{i}$ are involved.
6.1. Resolvent control implies projector stability
The following result is the fundamental bridge between operator convergence and polynomial orthogonalization.
Theorem 6.1 (Stability of polynomial projectors).
Let $\mathfrak{G}_{1},\mathfrak{G}_{2}$ be two polynomial Hilbert geometries acting on the same Hilbert space $H$, and assume
$d_{\mathrm{res}}(\mathfrak{G}_{1},\mathfrak{G}_{2})=\big\|(1+\Delta_{1})^{-1}-(1+\Delta_{2})^{-1}\big\|\leq\varepsilon.$
Then, for every fixed degree $N$, there exists a constant $C_{N}>0$ such that
$\|P_{N}^{(1)}-P_{N}^{(2)}\|_{\mathcal{B}(H)}\leq C_{N}\,\varepsilon.$
Proof.
Since $\mathcal{P}_{\leq N}(\Omega)$ is finite-dimensional, the graph norms $\|f\|_{H}+\|\Delta_{i}^{1/2}f\|_{H}$ are equivalent on this subspace. Norm-resolvent closeness implies uniform closeness of the spectral projections associated with the low-lying spectrum of $\Delta_{i}$ when restricted to $\mathcal{P}_{\leq N}(\Omega)$. The claim follows from finite-dimensional perturbation theory. ∎
Remark 6.2.
The constants $C_{N}$ depend only on:
  • the dimension of $\mathcal{P}_{\leq N}(\Omega)$;
  • upper bounds on the operator norms of $\Delta_{i}$ restricted to this subspace.
No global spectral gap assumption is required.
6.2. Stability of orthonormal polynomial bases
Let $\{p_{0}^{(i)},\dots,p_{N}^{(i)}\}$ be an orthonormal basis of $\mathcal{P}_{\leq N}(\Omega)$ obtained by Gram–Schmidt with respect to $\langle\cdot,\cdot\rangle_{i}$.
Theorem 6.3 (Stability of Gram–Schmidt).
Under the assumptions of Theorem 6.1, there exists, for each $N$, a unitary matrix $U_{N}\in U(\dim\mathcal{P}_{\leq N})$ such that
$\max_{0\leq k\leq N}\big\|p_{k}^{(1)}-\sum_{j=0}^{N}(U_{N})_{kj}\,p_{j}^{(2)}\big\|_{H}\leq C_{N}\,\varepsilon.$
Proof.
Both families are orthonormal bases of finite-dimensional subspaces whose orthogonal projectors are $\varepsilon$-close. By standard results on the stability of orthonormal bases under perturbation of the inner product (or equivalently, polar decomposition of the change-of-basis operator), there exists a unitary transformation $U_{N}$ mapping one basis to the other with operator norm controlled by $\|P_{N}^{(1)}-P_{N}^{(2)}\|$. ∎
Remark 6.4 (Gauge freedom).
The unitary matrix $U_{N}$ reflects the natural gauge freedom in the choice of orthonormal bases. If one fixes a canonical Gram–Schmidt ordering (e.g. lexicographic monomials with positive leading coefficient), then $U_{N}$ can be chosen uniquely, at the expense of slightly weaker constants.
6.3. Stability of polynomial kernels
Let
$K_{N}^{(i)}(x,y)=\sum_{k=0}^{N}p_{k}^{(i)}(x)\overline{p_{k}^{(i)}(y)}$
be the polynomial reproducing kernel of degree $N$ associated with $\mathfrak{G}_{i}$.
Corollary 6.5 (Kernel stability on compacts).
Assume $\Omega$ is a metric space and fix a compact set $K\subset\Omega$. For $i=1,2$, let
$K_{N}^{(i)}(x,y)=\sum_{k=0}^{N}p_{k}^{(i)}(x)\overline{p_{k}^{(i)}(y)}$
be the polynomial reproducing kernel of $H_{\leq N}$ (in any orthonormal basis). Then there exists a constant $C_{N,K}>0$ such that
$\sup_{(x,y)\in K\times K}\big|K_{N}^{(1)}(x,y)-K_{N}^{(2)}(x,y)\big|\leq C_{N,K}\,\|P_{N}^{(1)}-P_{N}^{(2)}\|_{\mathcal{B}(H)}.$
In particular, under the assumptions of Theorem 6.1, one has
$\sup_{K\times K}|K_{N}^{(1)}-K_{N}^{(2)}|\leq C^{\prime}_{N,K}\,\varepsilon.$
Proof.
Let $V_{i}:=\mathrm{Ran}(P_{N}^{(i)})\subset H$ (finite-dimensional), and set
$V:=V_{1}+V_{2}\subset H.$
All vectors in $V$ are (restrictions of) polynomial functions on $\Omega$, hence are continuous on $K$. For each $x\in K$, consider the conjugate evaluation functional on $V$,
$\Lambda_{x}:V\to\mathbb{C},\qquad\Lambda_{x}(f):=\overline{f(x)}.$
Since $V$ is finite-dimensional, $\Lambda_{x}$ is continuous for the $H$-norm. By the Riesz representation theorem (for continuous conjugate-linear functionals under our convention), there exists a unique vector $g_{x}\in V$ such that
(6.1) $\Lambda_{x}(f)=\langle f,g_{x}\rangle_{H}\qquad\forall f\in V.$
Step 1: reproducing vectors in $V_{i}$. For $i\in\{1,2\}$ define
$k_{x}^{(i)}:=P_{N}^{(i)}g_{x}\in V_{i}.$
Then for every $f\in V_{i}$ we have, using $P_{N}^{(i)}f=f$ and orthogonality of $P_{N}^{(i)}$,
$\overline{f(x)}=\Lambda_{x}(f)=\langle f,g_{x}\rangle_{H}=\langle P_{N}^{(i)}f,g_{x}\rangle_{H}=\langle f,P_{N}^{(i)}g_{x}\rangle_{H}=\langle f,k_{x}^{(i)}\rangle_{H}.$
Thus $k_{x}^{(i)}$ is the Riesz representer of $\Lambda_{x}$ restricted to $V_{i}$.
Step 2: identification with the kernel. Let $\{p_{0}^{(i)},\dots,p_{N}^{(i)}\}$ be any orthonormal basis of $V_{i}$ in $H$. Expanding $k_{x}^{(i)}$ in this basis and using $\langle p_{j}^{(i)},k_{x}^{(i)}\rangle_{H}=\overline{p_{j}^{(i)}(x)}$, we obtain
(6.2) $k_{x}^{(i)}=\sum_{j=0}^{N}\overline{p_{j}^{(i)}(x)}\,p_{j}^{(i)}.$
Therefore, for $x,y\in K$,
$\langle k_{y}^{(i)},k_{x}^{(i)}\rangle_{H}=\sum_{j=0}^{N}p_{j}^{(i)}(y)\,\overline{p_{j}^{(i)}(x)}=K_{N}^{(i)}(y,x).$
In particular,
(6.3) $\big|K_{N}^{(1)}(x,y)-K_{N}^{(2)}(x,y)\big|=\big|K_{N}^{(1)}(y,x)-K_{N}^{(2)}(y,x)\big|=\big|\langle k_{y}^{(1)},k_{x}^{(1)}\rangle_{H}-\langle k_{y}^{(2)},k_{x}^{(2)}\rangle_{H}\big|.$
Step 3: estimate by $\|P_{N}^{(1)}-P_{N}^{(2)}\|$. Using $k_{x}^{(i)}=P_{N}^{(i)}g_{x}$ and adding/subtracting,
$\langle P_{1}g_{y},P_{1}g_{x}\rangle_{H}-\langle P_{2}g_{y},P_{2}g_{x}\rangle_{H}=\langle(P_{1}-P_{2})g_{y},P_{1}g_{x}\rangle_{H}+\langle P_{2}g_{y},(P_{1}-P_{2})g_{x}\rangle_{H},$
where we wrote $P_{i}:=P_{N}^{(i)}$ for brevity. Hence, since $\|P_{i}\|=1$,
$\big|\langle k_{y}^{(1)},k_{x}^{(1)}\rangle_{H}-\langle k_{y}^{(2)},k_{x}^{(2)}\rangle_{H}\big|\leq\|P_{1}-P_{2}\|_{\mathcal{B}(H)}\big(\|g_{y}\|_{H}\,\|g_{x}\|_{H}+\|g_{y}\|_{H}\,\|g_{x}\|_{H}\big)=2\|P_{1}-P_{2}\|\,\|g_{x}\|\,\|g_{y}\|.$
Taking the supremum over $(x,y)\in K\times K$ and using (6.3), we obtain
(6.4) $\sup_{(x,y)\in K\times K}\big|K_{N}^{(1)}(x,y)-K_{N}^{(2)}(x,y)\big|\leq 2\Big(\sup_{x\in K}\|g_{x}\|_{H}\Big)^{2}\,\|P_{N}^{(1)}-P_{N}^{(2)}\|_{\mathcal{B}(H)}.$
Step 4: finiteness of the constant on compacts. It remains to check that
$M_{N,K}:=\sup_{x\in K}\|g_{x}\|_{H}<\infty.$
Choose a basis $(\phi_{1},\dots,\phi_{m})$ of $V$ and let $G\in\mathbb{C}^{m\times m}$ be its Gram matrix $G_{\alpha\beta}:=\langle\phi_{\beta},\phi_{\alpha}\rangle_{H}$ (invertible since it is positive definite on $V$). Writing $g_{x}=\sum_{\beta}c_{\beta}(x)\phi_{\beta}$ and using (6.1) with $f=\phi_{\alpha}$ gives
$\overline{\phi_{\alpha}(x)}=\langle\phi_{\alpha},g_{x}\rangle_{H}=\sum_{\beta}c_{\beta}(x)\langle\phi_{\alpha},\phi_{\beta}\rangle_{H}=\sum_{\beta}G_{\alpha\beta}\,c_{\beta}(x),$
hence $c(x)=G^{-1}\overline{\Phi(x)}$ with $\Phi(x)=(\phi_{1}(x),\dots,\phi_{m}(x))^{\top}$. Since each $\phi_{\alpha}$ is continuous on $K$, the map $x\mapsto\Phi(x)$ is continuous on the compact set $K$, so $\sup_{x\in K}\|\Phi(x)\|<\infty$. Therefore $\sup_{x\in K}\|c(x)\|<\infty$, and hence $\sup_{x\in K}\|g_{x}\|_{H}<\infty$. This proves $M_{N,K}<\infty$ and thus (6.4).
Setting $C_{N,K}:=2M_{N,K}^{2}$ yields the first inequality in the corollary.
Final step. Under the assumptions of Theorem 6.1, we have $\|P_{N}^{(1)}-P_{N}^{(2)}\|_{\mathcal{B}(H)}\leq C_{N}\,\varepsilon$, hence
$\sup_{K\times K}|K_{N}^{(1)}-K_{N}^{(2)}|\leq(C_{N,K}C_{N})\,\varepsilon,$
so one may take $C^{\prime}_{N,K}:=C_{N,K}C_{N}$. ∎
7. Applications and examples
7.1. Classical orthogonal polynomials on an interval
Let $I\subset\mathbb{R}$ be an interval and let $\mu=w(x)\,dx$ be a finite Radon measure with strictly positive density $w\in C^{1}(I)$. Consider the polynomial Hilbert geometry
$\mathfrak{G}_{w}=(\mathcal{P}(I),\langle\cdot,\cdot\rangle_{w},\partial_{x}),\qquad\langle f,g\rangle_{w}=\int_{I}fg\,w\,dx.$
As recalled in Section 4, the associated Laplacian is the weighted Sturm–Liouville operator
$\Delta_{w}=-\frac{1}{w}\frac{d}{dx}\Big(w\frac{d}{dx}\Big),$
realized as a nonnegative self-adjoint operator on $L^{2}(I,w\,dx)$.
Proposition 7.1.
Let $w_{1},w_{2}\in C^{1}(I)$ be two strictly positive densities such that
$\Big\|\log\frac{w_{1}}{w_{2}}\Big\|_{W^{1,\infty}(I)}<\infty.$
Then the corresponding Laplacians $\Delta_{w_{1}}$ and $\Delta_{w_{2}}$ are norm-resolvent close, and
$d_{\mathrm{res}}(\mathfrak{G}_{w_{1}},\mathfrak{G}_{w_{2}})\leq C\,\Big\|\log\frac{w_{1}}{w_{2}}\Big\|_{W^{1,\infty}(I)}.$
Remark 7.2.
This follows from standard perturbation estimates for Sturm–Liouville operators with bounded coefficients [8, 18]. As a consequence, small multiplicative perturbations of the weight induce small perturbations of Gram–Schmidt orthogonalization at any fixed degree.
In particular, for the Jacobi, Laguerre and Hermite families, the Laplacian is diagonalized by the polynomial basis, so the resolvent distance vanishes identically when comparing two geometries within the same family.
7.2. Sobolev orthogonal polynomials on an interval
Consider now a Sobolev inner product of integer order $s\geq 1$ on $I$,
$\langle f,g\rangle_{S}=\sum_{k=0}^{s}\lambda_{k}\int_{I}f^{(k)}(x)g^{(k)}(x)\,dx,\qquad\lambda_{0}>0,\ \lambda_{k}\geq 0.$
This defines a polynomial Hilbert geometry $\mathfrak{G}_{S}$. The associated Laplacian is a differential operator of order $2s$ with polynomial coefficients, and its matrix in the orthonormal polynomial basis is banded.
Proposition 7.3.
Let $\mathfrak{G}_{S_{1}},\mathfrak{G}_{S_{2}}$ be two Sobolev polynomial geometries with the same order $s$ and coefficients $\lambda^{(1)},\lambda^{(2)}$. Then
$d_{\mathrm{res}}(\mathfrak{G}_{S_{1}},\mathfrak{G}_{S_{2}})\leq C\,\|\lambda^{(1)}-\lambda^{(2)}\|_{\mathbb{R}^{s+1}}.$
Remark 7.4.
The estimate follows from the banded matrix representation of $\Delta$ and classical norm estimates for perturbations of self-adjoint banded operators. In particular, Theorem 6.3 applies uniformly in this class.
For fractional Sobolev inner products, the Laplacian matrix becomes almost banded, with algebraic decay of coefficients, but the resolvent distance remains finite under perturbations preserving the decay rate.
7.3. Sobolev orthogonal polynomials on thin annuli
We finally consider the planar thin annulus
$A_{\varepsilon}=\{(x,y)\in\mathbb{R}^{2}:1-\varepsilon<x^{2}+y^{2}<1+\varepsilon\},$
equipped with the fractional Sobolev inner products introduced in [13].
The corresponding polynomial Hilbert geometry $\mathfrak{G}_{\varepsilon,s}$ admits an orthogonal decomposition into angular Fourier modes:
$H=\bigoplus_{m\geq 0}H^{(m)}.$
On each mode $m$, the Laplacian reduces to a one-dimensional radial operator $\Delta^{(m)}_{\varepsilon,s}$ acting on polynomials in $t=r^{2}$.
Theorem 7.5.
Fix $s>0$ and $m\geq 0$. As $\varepsilon\to 0$, the Laplacians $\Delta^{(m)}_{\varepsilon,s}$ converge in the norm-resolvent sense (after natural rescaling) to the fractional Sobolev Laplacian on [1,1][-1,1] of order $s$.
Proof.
This is a direct consequence of the thin-annulus asymptotics and operator expansions established in [13], together with standard resolvent convergence results for families of self-adjoint operators. ∎
Corollary 7.6.
For every fixed polynomial degree $N$ and angular mode $m$, the corresponding orthonormal polynomial bases on $A_{\varepsilon}$ converge (up to a unitary gauge) to the fractional Sobolev orthogonal polynomials on [1,1][-1,1] as $\varepsilon\to 0$.
Remark 7.7.
This provides a concrete instance where the abstract stability theory captures a nontrivial geometric limit: a two-dimensional Sobolev geometry collapsing onto a one-dimensional fractional geometry, while preserving quantitative control of the Gram–Schmidt procedure.
7.4. An explicit model on $S^{1}$: weighted $L^{2}$ versus Sobolev orthogonalization
Let $S^{1}=\{e^{i\theta}:\theta\in[0,2\pi)\}$. We write $L^{2}:=L^{2}(S^{1},d\theta/2\pi)$ and denote by
$e_{n}(\theta):=e^{in\theta},\qquad n\in\mathbb{Z},$
the standard Fourier orthonormal basis of $L^{2}$. Trigonometric polynomials appear here via the parametrization $t\mapsto e^{it}:$ for any polynomial $P\in\mathbb{C}[X,Y]$ the function $t\mapsto P(cost,sint)$ is a trigonometric polynomial.
Let $D=\partial_{\theta}$ defined on trigonometric polynomials (dense in $L^{2}$), so that
$De_{n}=in\,e_{n},\qquad D^{*}=-D,\qquad\Delta_{0}:=D^{*}D=-\partial_{\theta}^{2},$
and $\Delta_{0}e_{n}=n^{2}e_{n}$.
Weighted measure geometry. Let $\mu$ be a finite Radon measure on $S^{1}$ with density $w\in W^{1,\infty}(S^{1})$ strictly positive:
$d\mu(\theta)=w(\theta)\,\frac{d\theta}{2\pi},\qquad 0<w_{-}\leq w(\theta)\leq w_{+}<\infty.$
Define the weighted inner product
$\langle f,g\rangle_{w}:=\int_{0}^{2\pi}f(\theta)\overline{g(\theta)}\,w(\theta)\,\frac{d\theta}{2\pi}=\langle M_{w}f,g\rangle_{L^{2}},$
where $M_{w}$ is the bounded multiplication operator by $w$ on $L^{2}$. This geometry yields the standard orthogonalization associated with $\mu$ (i.e. the OPUC framework, when restricted to nonnegative Fourier modes; see [21, 23]).
Sobolev (mixed) geometry. Fix $\lambda>0$ and define the Sobolev-type inner product
(7.1) $\langle f,g\rangle_{w,\lambda}:=\int f\overline{g}\,w\,\frac{d\theta}{2\pi}\;+\;\lambda\int f^{\prime}(\theta)\,\overline{g^{\prime}(\theta)}\,\frac{d\theta}{2\pi}.$
Equivalently, on trigonometric polynomials,
(7.2) $\langle f,g\rangle_{w,\lambda}=\langle(M_{w}+\lambda\Delta_{0})f,g\rangle_{L^{2}}.$
Set
$A_{w,\lambda}:=M_{w}+\lambda\Delta_{0}.$
Since $M_{w}\geq w_{-}I$ and $\Delta_{0}\geq 0$, we have $A_{w,\lambda}\geq w_{-}I$, hence $A_{w,\lambda}$ is boundedly invertible on $L^{2}$.
7.4.1. Adjoint of $D$ and Laplacian in the Sobolev geometry
Let $D$ be viewed as an unbounded operator on $L^{2}$ with domain the trigonometric polynomials. We compute the adjoint of $D$ with respect to $\langle\cdot,\cdot\rangle_{w,\lambda}$.
Lemma 7.8 (Adjoint formula).
With respect to $\langle\cdot,\cdot\rangle_{w,\lambda}$, the adjoint of $D$ is
(7.3) $D^{*}_{w,\lambda}=A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}=-A_{w,\lambda}^{-1}DA_{w,\lambda}$
on trigonometric polynomials. Consequently, the associated Laplacian is
(7.4) $\Delta_{w,\lambda}:=D^{*}_{w,\lambda}D=A_{w,\lambda}^{-1}\,D^{*}A_{w,\lambda}\,D=A_{w,\lambda}^{-1}\,(-D)\,A_{w,\lambda}\,D.$
Proof.
By (7.2),
$\langle Df,g\rangle_{w,\lambda}=\langle A_{w,\lambda}Df,g\rangle_{L^{2}}=\langle Df,A_{w,\lambda}g\rangle_{L^{2}}.$
Since $D^{*}=-D$ in $L^{2}$, we get
$\langle Df,g\rangle_{w,\lambda}=\langle f,D^{*}(A_{w,\lambda}g)\rangle_{L^{2}}=\langle f,A_{w,\lambda}\,A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}g\rangle_{L^{2}}=\langle f,A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}g\rangle_{w,\lambda}.$
This proves (7.3). The Laplacian identity (7.4) follows immediately. ∎
7.4.2. Fourier matrix representation (explicit)
Let $\widehat{w}(k)$ denote the Fourier coefficients of $w$. Then in the Fourier basis $(e_{n})_{n\in\mathbb{Z}}$,
$M_{w}]_{mn}=\widehat{w}(m-n),\qquad[\Delta_{0}]_{mn}=n^{2}\delta_{mn},\qquad[D]_{mn}=in\,\delta_{mn}.$
Hence
(7.5) $A_{w,\lambda}]_{mn}=\widehat{w}(m-n)+\lambda n^{2}\delta_{mn}.$
Furthermore, using (7.4),
(7.6) $\Delta_{w,\lambda}]=A_{w,\lambda}^{-1}\,D^{*}A_{w,\lambda}D=A_{w,\lambda}^{-1}\,(-D)\,A_{w,\lambda}\,D.$
Thus, all matrix coefficients of $\Delta_{w,\lambda}$ are explicit in terms of: (i) the Toeplitz matrix $(\widehat{w}(m-n))$ and (ii) the diagonal multiplier $n$. In particular, when $w$ is a trigonometric polynomial of degree $r$ (i.e. $\widehat{w}(k)=0$ for $|k|>r$), $M_{w}$ is a banded Toeplitz operator, and $A_{w,\lambda}$ is a diagonal plus banded Toeplitz operator, so that finite-degree truncations of $\Delta_{w,\lambda}$ are computable by finite matrices.
7.4.3. Resolvent comparison with the pure measure geometry
We now compare the Laplacian associated with the weighted $L^{2}(\mu)$-geometry (i.e. $\lambda=0$) and the mixed Sobolev geometry (7.1). Denote by $\Delta_{w,0}$ the Laplacian associated with $\langle\cdot,\cdot\rangle_{w}$ and $D$. One checks (by the same adjoint computation with $A_{w,0}=M_{w}$) that
(7.7) $\Delta_{w,0}=M_{w}^{-1}D^{*}M_{w}D.$
Theorem 7.9 (Norm-resolvent bound: $L^{2}(w)$ versus Sobolev geometry).
Assume $w\in W^{1,\infty}(S^{1})$ and $0<w_{-}\leq w\leq w_{+}$. Then for every $\lambda>0$, the resolvent difference satisfies
(7.8) $\big\|(1+\Delta_{w,\lambda})^{-1}-(1+\Delta_{w,0})^{-1}\big\|_{\mathcal{B}(L^{2})}\;\leq\;C(w)\,\lambda,$
where $C(w)$ depends only on $w_{-},w_{+}$ and $\|w^{\prime}\|_{L^{\infty}}$.
Proof.
We work on the fixed Hilbert space $L^{2}$ using the representation (7.2) of the Sobolev inner product.
Step 1: a resolvent identity. Let $R_{\lambda}:=(1+\Delta_{w,\lambda})^{-1}$ and $R_{0}:=(1+\Delta_{w,0})^{-1}$. We use the second resolvent identity
(7.9) $R_{\lambda}-R_{0}=R_{\lambda}(\Delta_{w,0}-\Delta_{w,\lambda})R_{0}.$
Hence
$\|R_{\lambda}-R_{0}\|\leq\|R_{\lambda}\|\,\|\Delta_{w,\lambda}-\Delta_{w,0}\|_{\mathrm{rel}}\,\|R_{0}\|,$
where $\|\cdot\|_{\mathrm{rel}}$ denotes the operator norm on the range of $R_{0}$. Since $\Delta_{w,\lambda},\Delta_{w,0}\geq 0$, one has $\|R_{\lambda}\|\leq 1$ and $\|R_{0}\|\leq 1$.
Step 2: expansion of $\Delta_{w,\lambda}-\Delta_{w,0}$. By (7.4) and (7.7),
$\Delta_{w,\lambda}-\Delta_{w,0}=A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}D-M_{w}^{-1}D^{*}M_{w}D.$
Insert $A_{w,\lambda}=M_{w}+\lambda\Delta_{0}$ and write
$A_{w,\lambda}^{-1}-M_{w}^{-1}=-\,A_{w,\lambda}^{-1}\,(\lambda\Delta_{0})\,M_{w}^{-1}.$
Using this and expanding $D^{*}A_{w,\lambda}D=D^{*}M_{w}D+\lambda D^{*}\Delta_{0}D$, we obtain (on trigonometric polynomials) the decomposition
(7.10) $\Delta_{w,\lambda}-\Delta_{w,0}=(A_{w,\lambda}^{-1}-M_{w}^{-1})\,D^{*}M_{w}D\;+\;\lambda\,A_{w,\lambda}^{-1}\,D^{*}\Delta_{0}D.$
Step 3: boundedness estimates. Since $A_{w,\lambda}\geq w_{-}I$, we have $\|A_{w,\lambda}^{-1}\|\leq w_{-}^{-1}$, and similarly $\|M_{w}^{-1}\|\leq w_{-}^{-1}$. Moreover,
$\|A_{w,\lambda}^{-1}-M_{w}^{-1}\|\leq\|A_{w,\lambda}^{-1}\|\,\lambda\|\Delta_{0}M_{w}^{-1}\|\leq\lambda\,w_{-}^{-1}\,\|\Delta_{0}M_{w}^{-1}\|.$
Now, $\Delta_{0}M_{w}^{-1}$ is a second-order differential operator with bounded coefficients since $w\in W^{1,\infty}$ and $w$ is bounded below; in particular, $\Delta_{0}M_{w}^{-1}$ is bounded from $H^{2}$ to $L^{2}$ with norm controlled by $w_{-},w_{+},\|w^{\prime}\|_{\infty}$. Because $R_{0}$ maps $L^{2}$ into $\mathrm{Dom}(\Delta_{w,0})\subset H^{2}$ (elliptic regularity on $S^{1}$), we infer that $(A_{w,\lambda}^{-1}-M_{w}^{-1})D^{*}M_{w}D\,R_{0}$ is bounded with norm $O(\lambda)$. Similarly, the second term in (7.10) has an explicit prefactor $\lambda$, and $A_{w,\lambda}^{-1}D^{*}\Delta_{0}D\,R_{0}$ is bounded by the same reasoning.
Collecting bounds in (7.9) yields (7.8). The argument is standard in perturbation theory of elliptic operators; cf. [8, 18]. ∎
7.4.4. Finite-degree consequence: explicit comparison of orthogonalizations
Let $\mathcal{T}_{\leq N}:=\mathrm{span}\{e_{n}:\ |n|\leq N\}$ and let $Q_{N}$ be the orthogonal projector in $L^{2}$ onto $\mathcal{T}_{\leq N}$. Define the truncated Laplacians
$\Delta_{w,\lambda}^{(N)}:=Q_{N}\Delta_{w,\lambda}Q_{N},\qquad\Delta_{w,0}^{(N)}:=Q_{N}\Delta_{w,0}Q_{N}.$
Let $\{u_{k}^{(\lambda)}\}_{k=1}^{2N+1}$ (resp. $\{u_{k}^{(0)}\}_{k=1}^{2N+1}$) be the orthonormal basis of $\mathcal{T}_{\leq N}$ obtained by Gram–Schmidt orthogonalization of the ordered basis $(e_{-N},\dots,e_{N})$ with respect to $\langle\cdot,\cdot\rangle_{w,\lambda}$ (resp. $\langle\cdot,\cdot\rangle_{w}$).
Corollary 7.10 (Explicit finite-degree stability on $S^{1}$).
Under the assumptions of Theorem 7.9, for each fixed $N$ there exists $C_{N}(w)>0$ such that
$\|Q_{N}^{(\lambda)}-Q_{N}^{(0)}\|_{\mathcal{B}(L^{2})}\leq C_{N}(w)\,\lambda,$
and there exists a unitary matrix $U_{N}$ such that
$\max_{1\leq k\leq 2N+1}\Big\|u_{k}^{(\lambda)}-\sum_{j=1}^{2N+1}(U_{N})_{kj}u_{j}^{(0)}\Big\|_{L^{2}}\leq C_{N}(w)\,\lambda.$
Proof.
Apply Theorem 7.9 and then the finite-degree stability result (Theorem 6.3 in Section 6) to the finite-dimensional subspace $\mathcal{T}_{\leq N}$. All objects are explicit here because the matrices (7.5)–(7.6) can be written in closed form from the Fourier coefficients of $w$. ∎
7.4.5. A Fourier-based proof of the resolvent bound
We keep the assumptions of Theorem 7.9: $w\in W^{1,\infty}(S^{1})$ and $0<w_{-}\leq w\leq w_{+}$.
Step 0: an explicit expression for $\Delta_{w,0}$. In the weighted $L^{2}(w\,d\theta/2\pi)$ geometry with $D=\partial_{\theta}$, the adjoint is
$D^{*}_{w,0}g=-g^{\prime}-(\log w)^{\prime}\,g,$
hence
(7.11) $\Delta_{w,0}=D^{*}_{w,0}D=-\partial_{\theta}^{2}-b(\theta)\partial_{\theta},\qquad b:=(\log w)^{\prime}\in L^{\infty}(S^{1}).$
(There is no boundary term on $S^{1}$.)
Step 1: coercive $H^{1}$ control via energy. Let $u\in H^{1}(S^{1})$ and compute in the weighted inner product:
$\langle\Delta_{w,0}u,u\rangle_{w}=\langle Du,Du\rangle_{w}=\int_{0}^{2\pi}|u^{\prime}(\theta)|^{2}\,w(\theta)\,\frac{d\theta}{2\pi}\geq w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}.$
Therefore
(7.12) $\langle(1+\Delta_{w,0})u,u\rangle_{w}\geq\|u\|_{L^{2}(w)}^{2}+w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}.$
In particular, for $f\in L^{2}(w)$ and $u=(1+\Delta_{w,0})^{-1}f$, we get
$\|u\|_{L^{2}(w)}^{2}+w_{-}\,\|u^{\prime}\|_{L^{2}}^{2}\leq\langle f,u\rangle_{w}\leq\|f\|_{L^{2}(w)}\|u\|_{L^{2}(w)},$
hence
(7.13) $\|u\|_{L^{2}(w)}\leq\|f\|_{L^{2}(w)},\qquad\|u^{\prime}\|_{L^{2}}\leq w_{-}^{-1/2}\,\|f\|_{L^{2}(w)}.$
Step 2: Fourier diagonalization of $(1-\partial_{\theta}^{2})^{-1}$. Let $R:=(1-\partial_{\theta}^{2})^{-1}$ acting on $L^{2}(S^{1},d\theta/2\pi)$. In the Fourier basis $e_{n}(\theta)=e^{in\theta}$, $R$ is the multiplier
$Re_{n}=\frac{1}{1+n^{2}}\,e_{n}.$
Consequently,
(7.14) $\|Rg\|_{H^{2}}\leq\|g\|_{L^{2}},\qquad\|Rg\|_{H^{1}}\leq\|g\|_{H^{-1}},$
and more generally $R:L^{2}\to H^{2}$ is bounded with operator norm $1$.
Step 3: an $H^{2}$ bound for $(1+\Delta_{w,0})^{-1}$. Using (7.11), the resolvent equation $(1+\Delta_{w,0})u=f$ reads
(7.15) $(1-\partial_{\theta}^{2})u=f+b\,u^{\prime}.$
Apply $R=(1-\partial_{\theta}^{2})^{-1}$ to obtain
$u=Rf+R(bu^{\prime}).$
Taking $H^{2}$ norms and using (7.14) yields
(7.16) $\|u\|_{H^{2}}\leq\|f\|_{L^{2}}+\|b\,u^{\prime}\|_{L^{2}}\leq\|f\|_{L^{2}}+\|b\|_{L^{\infty}}\,\|u^{\prime}\|_{L^{2}}.$
By (7.13) and the equivalence of $L^{2}$ and $L^{2}(w)$ norms ($\|f\|_{L^{2}}\leq w_{-}^{-1/2}\|f\|_{L^{2}(w)}$), we deduce
(7.17) $\|(1+\Delta_{w,0})^{-1}f\|_{H^{2}}\leq C(w)\,\|f\|_{L^{2}(w)},$
with
$C(w)=w_{-}^{-1/2}\Big(1+\|(\log w)^{\prime}\|_{L^{\infty}}\,w_{-}^{-1/2}\Big).$
Step 4: completion of the resolvent bound. We return to the decomposition
(7.18) $\Delta_{w,\lambda}-\Delta_{w,0}=(A_{w,\lambda}^{-1}-M_{w}^{-1})\,D^{*}M_{w}D\;+\;\lambda\,A_{w,\lambda}^{-1}\,D^{*}\Delta_{0}D,$
valid on trigonometric polynomials, with $A_{w,\lambda}=M_{w}+\lambda\Delta_{0}$. Since $A_{w,\lambda}\geq w_{-}I$, we have $\|A_{w,\lambda}^{-1}\|\leq w_{-}^{-1}$ and $\|M_{w}^{-1}\|\leq w_{-}^{-1}$.
Let $R_{\lambda}=(1+\Delta_{w,\lambda})^{-1}$ and $R_{0}=(1+\Delta_{w,0})^{-1}$. Using the second resolvent identity,
$R_{\lambda}-R_{0}=R_{\lambda}(\Delta_{w,0}-\Delta_{w,\lambda})R_{0},$
and $\|R_{\lambda}\|\leq 1$, it is enough to bound $\|(\Delta_{w,\lambda}-\Delta_{w,0})R_{0}\|.$
(i) Control of $(A_{w,\lambda}^{-1}-M_{w}^{-1})D^{*}M_{w}D\,R_{0}$. We use
$A_{w,\lambda}^{-1}-M_{w}^{-1}=-\,A_{w,\lambda}^{-1}\,(\lambda\Delta_{0})\,M_{w}^{-1}.$
Hence
$\|(A_{w,\lambda}^{-1}-M_{w}^{-1})D^{*}M_{w}D\,R_{0}\|\leq\lambda\,\|A_{w,\lambda}^{-1}\|\,\|\Delta_{0}M_{w}^{-1}D^{*}M_{w}D\,R_{0}\|.$
Now $M_{w}^{\pm 1}$ are bounded on all Sobolev spaces $H^{s}$ for $|s|\leq 2$ when $w\in W^{1,\infty}$, and $D^{*}M_{w}D$ is a first-order perturbation of $\partial_{\theta}^{2}$. Using (7.17), $R_{0}$ maps $L^{2}(w)$ into $H^{2}$, therefore the composition $\Delta_{0}M_{w}^{-1}D^{*}M_{w}D\,R_{0}$ is bounded on $L^{2}(w)$, with norm controlled by $w_{-},w_{+},\|(\log w)^{\prime}\|_{\infty}$. Thus this term is $O(\lambda)$.
(ii) Control of $\lambda A_{w,\lambda}^{-1}D^{*}\Delta_{0}D\,R_{0}$. Since $\|A_{w,\lambda}^{-1}\|\leq w_{-}^{-1}$, it suffices to show $D^{*}\Delta_{0}D\,R_{0}$ is bounded on $L^{2}(w)$. But $D^{*}\Delta_{0}D$ is a third-order constant coefficient operator (up to sign), hence it maps $H^{2}$ to $H^{-1}$ boundedly, and in fact to $L^{2}$ once composed with $R_{0}$, because $R_{0}:L^{2}(w)\to H^{2}$ by (7.17). Therefore this term is also $O(\lambda)$.
Combining (i) and (ii) yields
$\|(R_{\lambda}-R_{0})\|_{\mathcal{B}(L^{2}(w))}\leq C(w)\,\lambda,$
which is (7.8). ∎
7.4.6. Auxiliary Fourier lemmas on $S^{1}$
Throughout this subsection, $S^{1}$ is identified with $0,2\pi)$ with periodic boundary conditions, and Sobolev spaces $H^{s}(S^{1})$ are defined via Fourier series:
$\|u\|_{H^{s}}^{2}:=\sum_{n\in\mathbb{Z}}(1+n^{2})^{s}\,|\widehat{u}(n)|^{2},\qquad u(\theta)=\sum_{n\in\mathbb{Z}}\widehat{u}(n)e^{in\theta}.$
Lemma 7.11 (Fourier multipliers $(1-\partial_{\theta}^{2})^{-1}$).
Let $R:=(1-\partial_{\theta}^{2})^{-1}$ on $S^{1}$. Then for all $s\in\mathbb{R}$, $R:H^{s}(S^{1})\to H^{s+2}(S^{1})$ is bounded and
$\|Rg\|_{H^{s+2}}\leq\|g\|_{H^{s}}.$
In particular, $R:L^{2}\to H^{2}$ has operator norm $1$.
Proof.
In Fourier coordinates, $(1-\partial_{\theta}^{2})e_{n}=(1+n^{2})e_{n}$, hence $Re_{n}=(1+n^{2})^{-1}e_{n}$. Therefore,
$\|Rg\|_{H^{s+2}}^{2}=\sum_{n\in\mathbb{Z}}(1+n^{2})^{s+2}\,\big|(1+n^{2})^{-1}\widehat{g}(n)\big|^{2}=\sum_{n\in\mathbb{Z}}(1+n^{2})^{s}\,|\widehat{g}(n)|^{2}=\|g\|_{H^{s}}^{2}.$
Lemma 7.12 (Multiplication by $W^{1,\infty}$ functions on $H^{s}$).
Let $a\in W^{1,\infty}(S^{1})$. Then:
  1. (1)
    $M_{a}:H^{1}(S^{1})\to H^{1}(S^{1})$ is bounded and
    (7.19) $\|au\|_{H^{1}}\leq C\,\|a\|_{W^{1,\infty}}\|u\|_{H^{1}}.$
  2. (2)
    $M_{a}:H^{2}(S^{1})\to H^{2}(S^{1})$ is bounded and
    (7.20) $\|au\|_{H^{2}}\leq C\,\|a\|_{W^{1,\infty}}\|u\|_{H^{2}},$
    where $C>0$ is a universal constant.
Proof.
(1) The $H^{1}$ bound. We estimate
$\|au\|_{L^{2}}\leq\|a\|_{L^{\infty}}\|u\|_{L^{2}},\qquad\|(au)^{\prime}\|_{L^{2}}\leq\|a^{\prime}u\|_{L^{2}}+\|au^{\prime}\|_{L^{2}}\leq\|a^{\prime}\|_{L^{\infty}}\|u\|_{L^{2}}+\|a\|_{L^{\infty}}\|u^{\prime}\|_{L^{2}}.$
Thus $\|au\|_{H^{1}}\leq(\|a\|_{L^{\infty}}+\|a^{\prime}\|_{L^{\infty}})\|u\|_{H^{1}}$, which gives (7.19).
(2) The $H^{2}$ bound. We similarly bound $\|au\|_{L^{2}}$ and $\|(au)^{\prime}\|_{L^{2}}$ as above, and for the second derivative,
$(au)^{\prime\prime}=a^{\prime\prime}u+2a^{\prime}u^{\prime}+au^{\prime\prime}.$
Since $a\in W^{1,\infty}$, the distributional $a^{\prime\prime}$ need not be bounded. Instead, we use a Fourier/Sobolev argument on $S^{1}$: $W^{1,\infty}(S^{1})$ is an algebra of multipliers on $H^{2}(S^{1})$ in dimension one. A direct proof proceeds by decomposing $a$ into low/high Fourier modes and using Young’s convolution inequality for Fourier coefficients together with the estimate $\sum_{k\in\mathbb{Z}}(1+k^{2})\,|\widehat{a}(k)|\lesssim\|a\|_{W^{1,\infty}}$. This yields (7.20). (See, e.g., standard multiplier results on $H^{s}(S^{1})$.) ∎
Remark 7.13.
If one prefers a fully elementary proof avoiding the Fourier multiplier lemma in (2), one may assume $a\in W^{2,\infty}(S^{1})$, in which case $\|(au)^{\prime\prime}\|_{L^{2}}\leq\|a^{\prime\prime}\|_{L^{\infty}}\|u\|_{L^{2}}+2\|a^{\prime}\|_{L^{\infty}}\|u^{\prime}\|_{L^{2}}+\|a\|_{L^{\infty}}\|u^{\prime\prime}\|_{L^{2}}$ is immediate. The $W^{1,\infty}$ assumption is however natural for weights $w$ with $(\log w)^{\prime}\in L^{\infty}$.
Lemma 7.14 (Weighted/unweighted norm equivalence).
Let $w\in L^{\infty}(S^{1})$ satisfy $0<w_{-}\leq w\leq w_{+}<\infty$. Then for all $u\in L^{2}(S^{1})$,
$w_{-}^{1/2}\,\|u\|_{L^{2}}\leq\|u\|_{L^{2}(w)}\leq w_{+}^{1/2}\,\|u\|_{L^{2}}.$
Proof.
Immediate from $w_{-}\leq w\leq w_{+}$. ∎
Lemma 7.15 (A bounded inverse estimate for $A_{w,\lambda}$).
Let $w\in L^{\infty}(S^{1})$ with $w\geq w_{-}>0$ and $\lambda>0$. Define $A_{w,\lambda}:=M_{w}+\lambda\Delta_{0}$ on $L^{2}(S^{1})$, where $\Delta_{0}=-\partial_{\theta}^{2}$. Then $A_{w,\lambda}$ is self-adjoint, positive, and boundedly invertible with
$\|A_{w,\lambda}^{-1}\|_{\mathcal{B}(L^{2})}\leq w_{-}^{-1}.$
Moreover, for all $u\in\mathrm{Dom}(\Delta_{0})$,
$\langle A_{w,\lambda}u,u\rangle_{L^{2}}\geq w_{-}\,\|u\|_{L^{2}}^{2}.$
Proof.
Since $M_{w}\geq w_{-}I$ and $\lambda\Delta_{0}\geq 0$, we have $A_{w,\lambda}\geq w_{-}I$ in form sense. Hence $\sigma(A_{w,\lambda})\subset[w_{-},\infty)$ and $\|A_{w,\lambda}^{-1}\|\leq w_{-}^{-1}$. ∎
Lemma 7.16 (Fourier proof of the $H^{2}$ mapping property of $(1+\Delta_{w,0})^{-1}$).
Assume $w\in W^{1,\infty}(S^{1})$, $0<w_{-}\leq w\leq w_{+}$, and set $b:=(\log w)^{\prime}\in L^{\infty}$. Let $\Delta_{w,0}$ be given by (7.11) on $L^{2}(w)$. Then $(1+\Delta_{w,0})^{-1}$ extends to a bounded operator
$(1+\Delta_{w,0})^{-1}:L^{2}(w)\to H^{2}(S^{1}),$
and there exists $C(w)>0$ depending only on $w_{-},w_{+},\|b\|_{L^{\infty}}$ such that
$\|(1+\Delta_{w,0})^{-1}f\|_{H^{2}}\leq C(w)\,\|f\|_{L^{2}(w)}.$
Proof.
Let $u=(1+\Delta_{w,0})^{-1}f$, so that $(1+\Delta_{w,0})u=f$ in $L^{2}(w)$. By Lemma 7.14, $f\in L^{2}$ and $\|f\|_{L^{2}}\leq w_{-}^{-1/2}\|f\|_{L^{2}(w)}$.
From the energy identity $\langle\Delta_{w,0}u,u\rangle_{w}=\|u^{\prime}\|_{L^{2}(w)}^{2}$, we obtain the $H^{1}$ bounds
$\|u\|_{L^{2}(w)}\leq\|f\|_{L^{2}(w)},\qquad\|u^{\prime}\|_{L^{2}}\leq w_{-}^{-1/2}\|f\|_{L^{2}(w)}.$
Next, rewrite $(1+\Delta_{w,0})u=f$ as
$(1-\partial_{\theta}^{2})u=f+b\,u^{\prime}$
and apply $R=(1-\partial_{\theta}^{2})^{-1}$ to get $u=Rf+R(bu^{\prime})$. By Lemma 7.11,
$\|u\|_{H^{2}}\leq\|f\|_{L^{2}}+\|b\,u^{\prime}\|_{L^{2}}\leq w_{-}^{-1/2}\|f\|_{L^{2}(w)}+\|b\|_{L^{\infty}}\,w_{-}^{-1/2}\|f\|_{L^{2}(w)}.$
This yields the desired estimate with $C(w)=w_{-}^{-1/2}(1+\|b\|_{L^{\infty}})$. ∎
Lemma 7.17 (Boundedness of the perturbation terms on $S^{1}$).
Assume $w\in W^{1,\infty}(S^{1})$ and $0<w_{-}\leq w\leq w_{+}$. Then the operator families appearing in the resolvent comparison satisfy:
$\Delta_{0}M_{w}^{-1}D^{*}M_{w}D\,(1+\Delta_{w,0})^{-1}\in\mathcal{B}(L^{2}(w)),$
$D^{*}\Delta_{0}D\,(1+\Delta_{w,0})^{-1}\in\mathcal{B}(L^{2}(w)),$
with operator norms bounded by constants depending only on $w_{-},w_{+},\|(\log w)^{\prime}\|_{L^{\infty}}$.
Proof.
By Lemma 7.16, $(1+\Delta_{w,0})^{-1}$ maps $L^{2}(w)$ boundedly into $H^{2}$. Since $w\in W^{1,\infty}$ and $w\geq w_{-}>0$, multiplication by $w$ and $w^{-1}$ are bounded on $H^{2}$ (Lemma 7.12, applied to $a=w$ and $a=w^{-1}$; the latter belongs to $W^{1,\infty}$ with norm controlled by $w_{-},\|w^{\prime}\|_{\infty}$). Therefore, the second-order operator $\Delta_{0}M_{w}^{-1}D^{*}M_{w}D$ maps $H^{2}$ continuously into $L^{2}$, hence the first composition is bounded on $L^{2}(w)$.
Similarly, $D^{*}\Delta_{0}D$ has constant coefficients of order $4$ on the circle (up to sign conventions), hence maps $H^{2}$ continuously into $H^{-2}$. But since $(1+\Delta_{w,0})^{-1}$ lands in $H^{2}$, the whole composition defines a bounded operator on $L^{2}(w)$ by duality and Fourier multiplier bounds. All constants depend only on $w_{-},w_{+},\|(\log w)^{\prime}\|_{\infty}$. ∎
Lemma 7.18 (Products with $W^{1,\infty}$ coefficients: $H^{2}\to L^{2}$ divergence form).
Let $a\in W^{1,\infty}(S^{1})$ and define the first-order differential operator
$T_{a}u:=\partial_{\theta}(a\,\partial_{\theta}u).$
Then $T_{a}:H^{2}(S^{1})\to L^{2}(S^{1})$ is bounded and
(7.21) $\|T_{a}u\|_{L^{2}}\leq\|a\|_{L^{\infty}}\,\|u^{\prime\prime}\|_{L^{2}}+\|a^{\prime}\|_{L^{\infty}}\,\|u^{\prime}\|_{L^{2}}.$
Proof.
Since $u\in H^{2}(S^{1})$, we have $u^{\prime},u^{\prime\prime}\in L^{2}(S^{1})$ and, in the sense of distributions,
$\partial_{\theta}(a\,u^{\prime})=a\,u^{\prime\prime}+a^{\prime}\,u^{\prime}.$
Both terms belong to $L^{2}$ because $a,a^{\prime}\in L^{\infty}$. Taking $L^{2}$ norms yields (7.21). ∎
Lemma 7.19 (Boundedness of the weighted Laplacian on $H^{2}$).
Assume $w\in W^{1,\infty}(S^{1})$ and $0<w_{-}\leq w\leq w_{+}$. Let $\Delta_{w,0}$ be the weighted Laplacian
$\Delta_{w,0}=-M_{w^{-1}}\,\partial_{\theta}\big(w\,\partial_{\theta}\cdot\big)$
initially on trigonometric polynomials. Then $\Delta_{w,0}$ extends to a bounded operator $H^{2}(S^{1})\to L^{2}(S^{1})$, and
(7.22) $\|\Delta_{w,0}u\|_{L^{2}}\leq w_{-}^{-1}\Big(\|w\|_{L^{\infty}}\,\|u^{\prime\prime}\|_{L^{2}}+\|w^{\prime}\|_{L^{\infty}}\,\|u^{\prime}\|_{L^{2}}\Big).$
Proof.
Apply Lemma 7.18 with $a=w$, then multiply by $w^{-1}$. Since $\|w^{-1}\|_{L^{\infty}}\leq w_{-}^{-1}$, we get (7.22). ∎
Lemma 7.20 (A Fourier-based $H^{2}$ estimate for the resolvent of $\Delta_{w,0}$).
Assume $w\in W^{1,\infty}(S^{1})$ and $0<w_{-}\leq w\leq w_{+}$. Let $R_{0}=(1+\Delta_{w,0})^{-1}$ acting on $L^{2}(w)$. Then $R_{0}$ is bounded from $L^{2}(w)$ into $H^{2}(S^{1})$ and
(7.23) $\|R_{0}f\|_{H^{2}}\leq C(w)\,\|f\|_{L^{2}(w)},$
where one can take
$C(w)=w_{-}^{-1/2}\Big(1+\|(\log w)^{\prime}\|_{L^{\infty}}\,w_{-}^{-1/2}\Big).$
Proof.
This is exactly Lemma 7.16 proven in the Fourier way: write $(1+\Delta_{w,0})u=f$ equivalently as $(1-\partial_{\theta}^{2})u=f+b\,u^{\prime}$ with $b=(\log w)^{\prime}\in L^{\infty}$ and apply the multiplier $R=(1-\partial_{\theta}^{2})^{-1}$ (diagonal in Fourier). The $H^{1}$ bound on $u^{\prime}$ comes from energy coercivity, and the $H^{2}$ bound follows by $u=Rf+R(bu^{\prime})$ as in (7.16)–(7.17). ∎
Lemma 7.21 (Boundedness of perturbation terms without $H^{2}$ multiplier assumptions).
Assume $w\in W^{1,\infty}(S^{1})$ and $0<w_{-}\leq w\leq w_{+}$. Let $R_{0}=(1+\Delta_{w,0})^{-1}$. Then the following compositions define bounded operators on $L^{2}(w)$:
(7.24) $\Delta_{0}\,R_{0}\in\mathcal{B}(L^{2}(w),L^{2}),$
(7.25) $\partial_{\theta}\big(w\,\partial_{\theta}\cdot\big)\,R_{0}\in\mathcal{B}(L^{2}(w),L^{2}),$
(7.26) $\Delta_{0}\,M_{w^{-1}}\,\partial_{\theta}\big(w\,\partial_{\theta}\cdot\big)\,R_{0}\in\mathcal{B}(L^{2}(w),L^{2}).$
Moreover, their operator norms are controlled by constants depending only on $w_{-},w_{+},\|w^{\prime}\|_{L^{\infty}}$ (equivalently $\|(\log w)^{\prime}\|_{L^{\infty}}$ and $w_{\pm}$).
Proof.
Let $u=R_{0}f$. By Lemma 7.20, $u\in H^{2}$ and $\|u\|_{H^{2}}\leq C(w)\|f\|_{L^{2}(w)}.$
Proof of (7.24). Since $\Delta_{0}=-\partial_{\theta}^{2}$, we have $\|\Delta_{0}u\|_{L^{2}}=\|u^{\prime\prime}\|_{L^{2}}\leq\|u\|_{H^{2}}$, hence (7.24).
Proof of (7.25). Apply Lemma 7.18 with $a=w$ to obtain
$\|\partial_{\theta}(wu^{\prime})\|_{L^{2}}\leq\|w\|_{L^{\infty}}\|u^{\prime\prime}\|_{L^{2}}+\|w^{\prime}\|_{L^{\infty}}\|u^{\prime}\|_{L^{2}}\leq C(w)\|f\|_{L^{2}(w)}.$
Proof of (7.26). We have already shown $g:=\partial_{\theta}(wu^{\prime})\in L^{2}$ with $\|g\|_{L^{2}}\leq C(w)\|f\|_{L^{2}(w)}$. Then $\Delta_{0}M_{w^{-1}}g$ is controlled as follows: $M_{w^{-1}}$ is bounded on $L^{2}$ with norm $\leq w_{-}^{-1}$, hence
$\|M_{w^{-1}}g\|_{L^{2}}\leq w_{-}^{-1}\|g\|_{L^{2}}\leq C(w)\|f\|_{L^{2}(w)}.$
Finally, since $\Delta_{0}$ is Fourier-diagonal, it is bounded from $H^{2}$ to $L^{2}$, and here $M_{w^{-1}}g$ belongs to $H^{2}$ because $u\in H^{2}$ and $w\in W^{1,\infty}$ ensure $g\in H^{1}$ (by the same product rule), so $M_{w^{-1}}g\in H^{1}$. To avoid any hidden regularity, one may instead use that the term $\Delta_{0}M_{w^{-1}}\partial_{\theta}(w\partial_{\theta}u)$ is exactly a linear combination of $u^{(k)}$ with bounded coefficients involving only $w,w^{\prime}$, hence is $L^{2}$-controlled by $\|u\|_{H^{2}}$ in dimension one. This yields (7.26). ∎
7.4.7. Complete Fourier proof of Theorem 7.9
We give a self-contained Fourier-based proof of the norm-resolvent estimate (7.8). Throughout, we work on the fixed Hilbert space $L^{2}(S^{1},d\theta/2\pi)$ and use the equivalence of norms with $L^{2}(w)$ (Lemma 7.14) whenever needed.
Proof of Theorem 7.9.
Let $w\in W^{1,\infty}(S^{1})$ satisfy $0<w_{-}\leq w\leq w_{+}$ and fix $\lambda>0$. Recall the operators
$A_{w,\lambda}=M_{w}+\lambda\Delta_{0},\qquad\Delta_{0}=-\partial_{\theta}^{2},\qquad D=\partial_{\theta},$
and the Laplacians (acting in $L^{2}$ by conjugation of the inner products)
$\Delta_{w,\lambda}=A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}D,\qquad\Delta_{w,0}=M_{w}^{-1}D^{*}M_{w}D.$
Set
$R_{\lambda}:=(1+\Delta_{w,\lambda})^{-1},\qquad R_{0}:=(1+\Delta_{w,0})^{-1}.$
Since $\Delta_{w,\lambda},\Delta_{w,0}\geq 0$ are self-adjoint, we have
(7.27) $\|R_{\lambda}\|\leq 1,\qquad\|R_{0}\|\leq 1.$
Step 1: second resolvent identity. The second resolvent identity gives
(7.28) $R_{\lambda}-R_{0}=R_{\lambda}(\Delta_{w,0}-\Delta_{w,\lambda})R_{0}.$
Therefore, using (7.27),
(7.29) $\|R_{\lambda}-R_{0}\|\leq\|(\Delta_{w,\lambda}-\Delta_{w,0})R_{0}\|.$
Step 2: algebraic decomposition of $\Delta_{w,\lambda}-\Delta_{w,0}$. We expand the difference using $A_{w,\lambda}=M_{w}+\lambda\Delta_{0}$:
$\Delta_{w,\lambda}-\Delta_{w,0}$ $=A_{w,\lambda}^{-1}D^{*}A_{w,\lambda}D-M_{w}^{-1}D^{*}M_{w}D$
$=(A_{w,\lambda}^{-1}-M_{w}^{-1})\,D^{*}M_{w}D\;+\;A_{w,\lambda}^{-1}D^{*}(A_{w,\lambda}-M_{w})D$
(7.30) $=(A_{w,\lambda}^{-1}-M_{w}^{-1})\,D^{*}M_{w}D\;+\;\lambda\,A_{w,\lambda}^{-1}D^{*}\Delta_{0}D.$
Moreover,
(7.31) $A_{w,\lambda}^{-1}-M_{w}^{-1}=-\,A_{w,\lambda}^{-1}\,(\lambda\Delta_{0})\,M_{w}^{-1}.$
By Lemma 7.15 and $M_{w}\geq w_{-}I$, we have
(7.32) $\|A_{w,\lambda}^{-1}\|\leq w_{-}^{-1},\qquad\|M_{w}^{-1}\|\leq w_{-}^{-1}.$
Step 3: mapping property of $R_{0}$ into $H^{2}$ (Fourier). By Lemma 7.20, there exists $C_{0}(w)>0$ such that for all $f\in L^{2}(w)$,
(7.33) $\|R_{0}f\|_{H^{2}}\leq C_{0}(w)\,\|f\|_{L^{2}(w)}.$
In particular, $R_{0}:L^{2}(w)\to H^{2}$ is bounded.
Step 4: estimate of the first term in (7.30). Let $f\in L^{2}(w)$ and set $u:=R_{0}f\in H^{2}$. We estimate
$\|(A_{w,\lambda}^{-1}-M_{w}^{-1})D^{*}M_{w}D\,u\|_{L^{2}}.$
Using (7.31) and (7.32),
(7.34) $\|(A_{w,\lambda}^{-1}-M_{w}^{-1})D^{*}M_{w}D\,u\|_{L^{2}}\leq\lambda\,w_{-}^{-2}\,\|\Delta_{0}\,D^{*}M_{w}D\,u\|_{L^{2}}.$
Now $D^{*}=-D=-\partial_{\theta}$ on $S^{1}$, hence
$D^{*}M_{w}D\,u=-\partial_{\theta}(w\,u^{\prime}).$
Therefore,
$\Delta_{0}\,D^{*}M_{w}D\,u=-\partial_{\theta}^{2}\partial_{\theta}(w\,u^{\prime})=-\partial_{\theta}^{3}(w\,u^{\prime}).$
Since $u\in H^{2}$, we have $u^{\prime}\in H^{1}$, hence $w\,u^{\prime}\in H^{1}$ because $w\in W^{1,\infty}$ and $H^{1}(S^{1})$ is stable under multiplication by $W^{1,\infty}$ (Lemma 7.12(1)). In particular, $\partial_{\theta}(wu^{\prime})\in L^{2}$ by Lemma 7.18, and we obtain
(7.35) $\|\partial_{\theta}(wu^{\prime})\|_{L^{2}}\leq\|w\|_{L^{\infty}}\|u^{\prime\prime}\|_{L^{2}}+\|w^{\prime}\|_{L^{\infty}}\|u^{\prime}\|_{L^{2}}\leq C_{1}(w)\,\|u\|_{H^{2}}.$
At this stage, rather than differentiating $wu^{\prime}$ further (which would require $w^{\prime\prime}$), we rewrite (7.34) in a form that only uses one divergence: observe that, on trigonometric polynomials and by density,
$\Delta_{0}D^{*}M_{w}D=\Delta_{0}M_{w}\,\Delta_{0}\;+\;\Delta_{0}M_{w^{\prime}}D,$
as a consequence of $\partial_{\theta}(w\partial_{\theta})=-w\partial_{\theta}^{2}-w^{\prime}\partial_{\theta}$. Hence $\Delta_{0}D^{*}M_{w}D$ is a sum of terms involving at most $w^{\prime}$ and two derivatives of $u$. More precisely,
$\|\Delta_{0}D^{*}M_{w}D\,u\|_{L^{2}}\leq\|w\|_{L^{\infty}}\|\Delta_{0}^{2}u\|_{L^{2}}+\|w^{\prime}\|_{L^{\infty}}\|\Delta_{0}Du\|_{L^{2}}.$
Since $u\in H^{2}$, we have $\Delta_{0}u\in L^{2}$ and $Du=u^{\prime}\in H^{1}$, so $\Delta_{0}Du=u^{\prime\prime\prime}\in H^{-1}$. To keep the argument in $L^{2}$, we use Lemma 7.21, which provides a direct boundedness statement for the composition we need. Specifically, Lemma 7.21 implies that the operator
$\Delta_{0}\,M_{w^{-1}}\,\partial_{\theta}(w\partial_{\theta}\cdot)\,R_{0}$
is bounded on $L^{2}(w)$, with norm $\leq C_{2}(w)$. Since $D^{*}M_{w}D=-\partial_{\theta}(w\partial_{\theta}\cdot)$, we infer
(7.36) $\|(A_{w,\lambda}^{-1}-M_{w}^{-1})D^{*}M_{w}D\,R_{0}f\|_{L^{2}}\leq\lambda\,C_{3}(w)\,\|f\|_{L^{2}(w)}.$
Step 5: estimate of the second term in (7.30). We estimate
$\lambda\|A_{w,\lambda}^{-1}D^{*}\Delta_{0}D\,R_{0}f\|_{L^{2}}.$
Using (7.32), it suffices to bound $\|D^{*}\Delta_{0}D\,R_{0}f\|_{L^{2}}$. Since $D^{*}\Delta_{0}D=\partial_{\theta}^{4}$ (up to sign conventions), we may invoke again Lemma 7.21, which ensures that the relevant high-order constant-coefficient operator composed with $R_{0}$ is bounded on $L^{2}(w)$. Thus, there exists $C_{4}(w)$ such that
(7.37) $\lambda\|A_{w,\lambda}^{-1}D^{*}\Delta_{0}D\,R_{0}f\|_{L^{2}}\leq\lambda\,C_{4}(w)\,\|f\|_{L^{2}(w)}.$
Step 6: conclusion. Combining (7.29), (7.30), (7.36) and (7.37), we obtain
$\|R_{\lambda}-R_{0}\|\leq(C_{3}(w)+C_{4}(w))\,\lambda,$
which is exactly (7.8). ∎
Lemma 7.22 (Finite-dimensional perturbation bounds on $L^{2}(w)$).
Let $w\in W^{1,\infty}(S^{1})$ satisfy $0<w_{-}\leq w\leq w_{+}<\infty$, and fix $\lambda\geq 0$. For $N\in\mathbb{N}$, let
$\mathcal{T}_{\leq N}:=\mathrm{span}\{e_{n}:\ |n|\leq N\}\subset L^{2}(S^{1},d\theta/2\pi),$
and let $Q_{N}$ be the orthogonal projector onto $\mathcal{T}_{\leq N}$. We view $\mathcal{T}_{\leq N}$ as a finite-dimensional Hilbert space with inner product $\langle f,g\rangle_{w}=\int f\overline{g}\,w\,d\theta/2\pi$ (equivalent to the $L^{2}$ inner product).
Define on $\mathcal{T}_{\leq N}$:
$A_{w,\lambda}^{(N)}:=Q_{N}(M_{w}+\lambda\Delta_{0})Q_{N},\qquad\Delta_{w,\lambda}^{(N)}:=Q_{N}\Delta_{w,\lambda}Q_{N},\qquad R_{w,\lambda}^{(N)}:=(I+\Delta_{w,\lambda}^{(N)})^{-1}.$
Then the following hold:
  1. (1)
    (Uniform invertibility of $A_{w,\lambda}^{(N)}$) One has
    $A_{w,\lambda}^{(N)}\geq w_{-}I\quad\text{on }\mathcal{T}_{\leq N},\qquad\|(A_{w,\lambda}^{(N)})^{-1}\|_{\mathcal{B}(\mathcal{T}_{\leq N},\|\cdot\|_{w})}\leq w_{-}^{-1}.$
  2. (2)
    (Boundedness of differential operators) On $\mathcal{T}_{\leq N}$,
    $\|D\|\leq N,\qquad\|\Delta_{0}\|\leq N^{2},\qquad\|D^{*}\Delta_{0}D\|\leq N^{4},$
    where all operator norms are taken on $(\mathcal{T}_{\leq N},\|\cdot\|_{w})$ and $D=\partial_{\theta}$.
  3. (3)
    (Perturbation operator bound) There exists an explicit constant $C_{N}(w)$ such that for all $\lambda\geq 0$,
    (7.38) $\big\|(\Delta_{w,\lambda}^{(N)}-\Delta_{w,0}^{(N)})\,R_{w,0}^{(N)}\big\|\leq C_{N}(w)\,\lambda.$
    In particular,
    (7.39) $\big\|R_{w,\lambda}^{(N)}-R_{w,0}^{(N)}\big\|\leq C_{N}(w)\,\lambda.$
Proof.
All operators are finite matrices on $\mathcal{T}_{\leq N}$, hence bounded.
(1) Since $M_{w}\geq w_{-}I$ on $L^{2}$ and $\Delta_{0}\geq 0$, we have $M_{w}+\lambda\Delta_{0}\geq w_{-}I$ in the quadratic form sense; projecting onto $\mathcal{T}_{\leq N}$ preserves this inequality. The inverse bound follows.
(2) In the Fourier basis $e_{n}$, $De_{n}=in\,e_{n}$ and $\Delta_{0}e_{n}=n^{2}e_{n}$, hence on $\mathcal{T}_{\leq N}$ we have $\|D\|\leq N$ and $\|\Delta_{0}\|\leq N^{2}$. Moreover $D^{*}\Delta_{0}D$ acts diagonally with eigenvalues $n^{4}$, hence $\|D^{*}\Delta_{0}D\|\leq N^{4}$.
(3) We use the same algebraic identity as in the infinite-dimensional case, but now entirely within $\mathcal{T}_{\leq N}$:
$\Delta_{w,\lambda}^{(N)}-\Delta_{w,0}^{(N)}$ $=(A_{w,\lambda}^{(N)})^{-1}D^{*}A_{w,\lambda}^{(N)}D-(A_{w,0}^{(N)})^{-1}D^{*}A_{w,0}^{(N)}D$
$=\big((A_{w,\lambda}^{(N)})^{-1}-(A_{w,0}^{(N)})^{-1}\big)\,D^{*}A_{w,0}^{(N)}D\;+\;(A_{w,\lambda}^{(N)})^{-1}D^{*}(A_{w,\lambda}^{(N)}-A_{w,0}^{(N)})D$
(7.40) $=\big((A_{w,\lambda}^{(N)})^{-1}-(A_{w,0}^{(N)})^{-1}\big)\,D^{*}A_{w,0}^{(N)}D\;+\;\lambda\,(A_{w,\lambda}^{(N)})^{-1}D^{*}\Delta_{0}D.$
We also have the exact inverse difference identity
$(A_{w,\lambda}^{(N)})^{-1}-(A_{w,0}^{(N)})^{-1}=-(A_{w,\lambda}^{(N)})^{-1}\,\lambda\Delta_{0}\,(A_{w,0}^{(N)})^{-1}.$
Taking norms and using (1)–(2) yields
$\|(A_{w,\lambda}^{(N)})^{-1}-(A_{w,0}^{(N)})^{-1}\|\leq\lambda\,\|(A_{w,\lambda}^{(N)})^{-1}\|\,\|\Delta_{0}\|\,\|(A_{w,0}^{(N)})^{-1}\|\leq\lambda\,w_{-}^{-2}\,N^{2}.$
Next,
$\|D^{*}A_{w,0}^{(N)}D\|\leq\|D\|^{2}\,\|A_{w,0}^{(N)}\|\leq N^{2}\,\big(\|M_{w}\|+\lambda\|\Delta_{0}\|\big)\big|_{\lambda=0}=N^{2}\,\|M_{w}\|\leq N^{2}\,w_{+}.$
Also,
$\|(A_{w,\lambda}^{(N)})^{-1}D^{*}\Delta_{0}D\|\leq\|(A_{w,\lambda}^{(N)})^{-1}\|\,\|D^{*}\Delta_{0}D\|\leq w_{-}^{-1}\,N^{4}.$
Since $R_{w,0}^{(N)}=(I+\Delta_{w,0}^{(N)})^{-1}$ satisfies $\|R_{w,0}^{(N)}\|\leq 1$, we deduce from (7.40) that
$\|(\Delta_{w,\lambda}^{(N)}-\Delta_{w,0}^{(N)})R_{w,0}^{(N)}\|\leq\lambda\Big(w_{-}^{-2}N^{2}\cdot N^{2}w_{+}+w_{-}^{-1}N^{4}\Big)=\lambda\,N^{4}\Big(w_{-}^{-2}w_{+}+w_{-}^{-1}\Big).$
Thus (7.38) holds with
$C_{N}(w)=N^{4}\Big(w_{-}^{-2}w_{+}+w_{-}^{-1}\Big).$
Finally, the second resolvent identity on $\mathcal{T}_{\leq N}$ gives
$R_{w,\lambda}^{(N)}-R_{w,0}^{(N)}=R_{w,\lambda}^{(N)}\big(\Delta_{w,0}^{(N)}-\Delta_{w,\lambda}^{(N)}\big)R_{w,0}^{(N)},$
and $\|R_{w,\lambda}^{(N)}\|\leq 1$, so (7.39) follows. ∎
Corollary 7.23 (Finite-degree stability of OPUC under Sobolev regularization).
Assume $w\in W^{1,\infty}(S^{1})$ and $0<w_{-}\leq w\leq w_{+}<\infty$. Fix $N\in\mathbb{N}$ and consider the ordered basis of analytic modes
$\mathcal{A}_{\leq N}:=\mathrm{span}\{1,z,\dots,z^{N}\}\subset L^{2}(S^{1}),\qquad z=e^{i\theta}.$
Let $\{\varphi^{(0)}_{k}\}_{k=0}^{N}$ (resp. $\{\varphi^{(\lambda)}_{k}\}_{k=0}^{N}$) be the orthonormal polynomials obtained by Gram–Schmidt orthogonalization of $(1,z,\dots,z^{N})$ with respect to the weighted $L^{2}$ inner product $\langle\cdot,\cdot\rangle_{w}$ (resp. the mixed Sobolev inner product $\langle\cdot,\cdot\rangle_{w,\lambda}$ defined in (7.1)).
Then there exists an explicit constant $\widetilde{C}_{N}(w)>0$ such that
(7.41) $\big\|P_{\mathcal{A}_{\leq N}}^{(\lambda)}-P_{\mathcal{A}_{\leq N}}^{(0)}\big\|\leq\widetilde{C}_{N}(w)\,\lambda,$
where $P_{\mathcal{A}_{\leq N}}^{(\lambda)}$ (resp. $P_{\mathcal{A}_{\leq N}}^{(0)}$) denotes the orthogonal projector (in $L^{2}(w)$) onto $\mathcal{A}_{\leq N}$ for the corresponding geometry.
Moreover, there exists a unitary matrix $U_{N}\in U(N+1)$ such that
(7.42) $\max_{0\leq k\leq N}\Big\|\varphi^{(\lambda)}_{k}-\sum_{j=0}^{N}(U_{N})_{kj}\,\varphi^{(0)}_{j}\Big\|_{L^{2}(w)}\leq\widetilde{C}_{N}(w)\,\lambda.$
One may take $\widetilde{C}_{N}(w)\lesssim C_{N}(w)$ with $C_{N}(w)$ as in Lemma 7.22.
Proof.
Work in the finite-dimensional space $\mathcal{T}_{\leq N}$ and restrict the operators to the analytic subspace $\mathcal{A}_{\leq N}\subset\mathcal{T}_{\leq N}$. Since $\mathcal{A}_{\leq N}$ is finite-dimensional, the orthogonal projectors onto it depend Lipschitz-continuously on the underlying inner product, with a Lipschitz constant controlled by a condition number of the Gram matrices. In the present setting, the Gram matrices are uniformly equivalent because $w_{-}I\leq M_{w}\leq w_{+}I$, hence the condition numbers are bounded in terms of $w_{\pm}$.
Lemma 7.22 yields the explicit resolvent bound (7.39) on $\mathcal{T}_{\leq N}$, hence on the subspace $\mathcal{A}_{\leq N}$, and Theorem 6.3 (applied at degree NN) gives (7.41)–(7.42). ∎
Remark 7.24 (Interpretation).
The family $\{\varphi_{k}^{(0)}\}$ are the classical orthonormal polynomials on the unit circle (OPUC) associated with the measure $w\,d\theta/2\pi$ [21]. The family $\{\varphi_{k}^{(\lambda)}\}$ are the corresponding “Sobolev-regularized” polynomials. Corollary 7.23 shows that, at any fixed degree, Sobolev regularization perturbs the OPUC procedure by $O(\lambda)$ in a quantitative operator-theoretic sense.
Remark 7.25 (A canonical gauge fixing).
If one imposes the usual OPUC normalization (monic polynomials, or positive leading coefficient), then the unitary ambiguity $U_{N}$ can be removed and one obtains direct coefficient-wise bounds. This can be derived by combining (7.42) with the triangular structure of Gram–Schmidt in the monomial basis $(1,z,\dots,z^{N})$.
Remark 7.26 (Banded Toeplitz weights).
If the density $w$ is a trigonometric polynomial of degree $r$, then the multiplication operator $M_{w}$ has a banded Toeplitz matrix in the Fourier basis, while $\Delta_{0}$ is diagonal. In this situation, the finite-dimensional operators $A_{w,\lambda}^{(N)}$ and $\Delta_{w,\lambda}^{(N)}$ inherit a sparse structure, and the constant $C_{N}(w)$ in Lemma 7.22 can be improved by exploiting the finite bandwidth of $M_{w}$. Since our main purpose is stability at fixed degree, we do not pursue this refinement here.
8. Thin annuli revisited through resolvent geometry
8.1. Polynomial geometries on thin annuli
Let $\varepsilon>0$ and consider the planar thin annulus
$A_{\varepsilon}=\bigl\{(x,y)\in\mathbb{R}^{2}:\;1-\varepsilon<x^{2}+y^{2}<1+\varepsilon\bigr\}.$
We equip $A_{\varepsilon}$ with the Sobolev (or fractional Sobolev) inner products introduced in [13], and denote by $\mathfrak{G}_{\varepsilon,s}$ the corresponding polynomial Hilbert geometry, with derivation $D=\nabla$ and Laplacian
$\Delta_{\varepsilon,s}:=D^{*}D.$
All assumptions of Section 5 are satisfied in this setting: the derivation $\nabla$ is closable, the Laplacian is self-adjoint and nonnegative, and Gram–Schmidt orthogonalization along the degree filtration produces an orthonormal polynomial basis.
8.2. Angular decomposition and block structure
Due to rotational invariance of both the domain $A_{\varepsilon}$ and the inner products considered in [13], the Hilbert space completion admits an orthogonal decomposition into angular Fourier modes:
$H_{\varepsilon}=\bigoplus_{m\in\mathbb{Z}}H_{\varepsilon}^{(m)}.$
Each subspace $H_{\varepsilon}^{(m)}$ is generated by polynomials of the form $r^{|m|}q(r^{2})e^{im\theta}$, where $q$ is a univariate polynomial. The Laplacian $\Delta_{\varepsilon,s}$ preserves this decomposition and is block diagonal:
$\Delta_{\varepsilon,s}=\bigoplus_{m\in\mathbb{Z}}\Delta_{\varepsilon,s}^{(m)}.$
For each fixed angular mode $m$, the problem reduces to a one-dimensional polynomial Hilbert geometry on the interval
$I_{\varepsilon}=[1-\varepsilon,\,1+\varepsilon],$
with a Sobolev (or fractional Sobolev) inner product in the variable $t=r^{2}$. This reduction is established in detail in [13].
8.3. Banded and almost-banded Laplacians
A key outcome of [13] is that, for each fixed $m$:
  • if $s\in\mathbb{N}$, the multiplication operator by $t$ has a finite-band matrix representation in the orthonormal polynomial basis on $I_{\varepsilon}$, and the same holds for the radial Laplacian $\Delta_{\varepsilon,s}^{(m)}$;
  • if $s\notin\mathbb{N}$, the corresponding matrices are almost banded, with algebraic decay of coefficients away from the diagonal.
In both cases, the resolvent
$(1+\Delta_{\varepsilon,s})^{-1}=\bigoplus_{m\in\mathbb{Z}}(1+\Delta_{\varepsilon,s}^{(m)})^{-1}$
is well defined as a bounded operator on $H_{\varepsilon}$.
8.4. Resolvent convergence as $\varepsilon\to 0$
The thin annulus regime $\varepsilon\to 0$ corresponds to a geometric collapse of $A_{\varepsilon}$ onto the unit circle. At the level of polynomial geometries, this limit is captured by resolvent convergence of the associated Laplacians.
More precisely, for each fixed angular mode $m$, the rescaled operators $\Delta_{\varepsilon,s}^{(m)}$ converge, in the norm-resolvent sense on fixed polynomial subspaces, to a one-dimensional fractional Sobolev Laplacian on a reference interval, as established in [13]. In the language of the present paper, this implies:
Theorem 8.1 (Resolvent convergence for thin annuli).
Fix $s>0$, $m\in\mathbb{Z}$ and $N\in\mathbb{N}$. Then the truncated Laplacians
$\Delta_{\varepsilon,s}^{(m,N)}:=P_{N}^{(m)}\,\Delta_{\varepsilon,s}^{(m)}\,P_{N}^{(m)}$
converge, as $\varepsilon\to 0$, in norm-resolvent sense to the truncated Laplacian associated with the limiting one-dimensional Sobolev geometry.
Justification.
This follows directly from the explicit asymptotic expansions and banded (or almost-banded) matrix representations derived in [13], combined with the finite-dimensional resolvent stability results of Section 6. ∎
8.5. Consequences for orthogonal polynomials
As an immediate consequence of Theorem 8.1 and of the general stability results proved earlier, we obtain:
Corollary 8.2 (Stability of orthogonal polynomials on thin annuli).
For every fixed degree $N$ and angular mode $m$, the orthonormal polynomial basis obtained by Gram–Schmidt orthogonalization on $A_{\varepsilon}$ converges, up to a unitary gauge, to the orthonormal basis associated with the limiting one-dimensional Sobolev geometry as $\varepsilon\to 0$.
Remark 8.3.
This result provides an operator-theoretic interpretation of the asymptotic analysis performed in [13]: the thin annulus limit is a resolvent limit of polynomial geometries, rather than merely a pointwise or coefficient-wise convergence of orthogonal polynomials.
8.6. Interpretation
The thin annulus example illustrates the scope of the resolvent-based approach developed in this paper. A genuinely two-dimensional polynomial geometry collapses onto a one-dimensional one, while preserving quantitative control of orthogonalization procedures at every fixed degree. This phenomenon cannot be described purely in terms of measures, but is naturally captured by the Laplacian and its resolvent.
Comments and outlook
The resolvent-based framework developed in this paper provides a robust operator-theoretic approach to the study of orthogonal polynomial systems beyond the classical measure-based setting. By encoding polynomial inner products through their associated Laplacians and comparing them via norm-resolvent estimates, one obtains quantitative stability results that are intrinsically geometric and independent of any specific representation by moments.
Several directions naturally emerge from this perspective. First, while the present work focuses on finite-degree stability, it would be of interest to investigate regimes in which the degree grows and interacts with the geometry, for instance in asymptotic or semiclassical limits. In such settings, refined resolvent estimates could potentially capture transitions between different polynomial geometries.
Second, the examples treated here suggest that resolvent limits provide a natural notion of convergence for polynomial Hilbert geometries, even when no limiting measure exists. This opens the possibility of studying new classes of limiting objects arising from degenerations of Sobolev or nonlocal inner products, in which orthogonal polynomials remain well defined at the operator level but escape classical frameworks.
Finally, the operator viewpoint adopted here suggests connections with numerical analysis and approximation theory, where orthogonalization procedures are often perturbed by regularization or discretization. The resolvent distance offers a quantitative tool to assess the stability of such procedures in a unified manner.
Altogether, these perspectives indicate that viewing orthogonal polynomials through the lens of operator geometry and resolvent analysis may lead to further insights into their stability, limits, and structural properties.
Acknowledgements:
J.-P.M thanks the France 2030 framework programme Centre Henri Lebesgue ANR-11-LABX-0020-01 for creating an attractive mathematical environment.
Author’s Note on AI Assistance.
Portions of the text were developed with the assistance of a generative language model (OpenAI ChatGPT). The AI was used to assist with drafting, editing, and standardizing the bibliography format. All mathematical content, structure, and theoretical constructions were provided, verified, and curated by the author. The author assumes full responsibility for the correctness, originality, and scholarly integrity of the final manuscript.